### v2

# LMM-Assisted Breast Cancer Treatment Target Segmentation with Consistency Embedding

Abstract
Recent advancements in Artificial Intelligence (AI) have profoundly influenced medical fields, by providing tools to reduce clinical workloads. However, most AI models are constrained to execute unimodal tasks, in stark contrast to the comprehensive approaches utilized by medical professionals. To address this, here we present RO-LMM, a multi-purpose large multimodal model (LMM) tailored for the field of radiation oncology. This model covers series of tasks within clinical workflow, adept at clinical report summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation. In particular, to perform consecutive clinical tasks, we further present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the capability of handling clean inputs, and transform this concept into LMM-driven segmentation framework as Consistency Embedding Segmentation (CESEG). Experimental results on multi-centre cohorts demonstrate our RO-LMM’s promising performance for multiple clinical tasks with generalization capabilities.
Figure 1: RO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology. The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation. All the patient information and any potential personally identifiable information are de-identifiied, and the clinical data usage is approved by IRB.
**footnotetext: These authors contributed equally to this work\$$\dagger$\$$\dagger$\$footnotetext: Corresponding authors
1 Introduction
Recently, the emergence of a new generation of AI models, known as foundation models, marks a significant departure from previous paradigms [34]. These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [3, 10, 12, 27, 31, 25].
These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision-making. Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [21]. Specifically, in contrast to these unimodal AIs, generalist medical AIs can encompass a holistic understanding of clinical workflows, which can receive and reason a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [41, 46, 39, 50, 34].
Inspired by this paradigm shift in medical AIs, here we present RO-LMM, a prototype of a multi-purpose medical large multimodal model (LMM), tailored to support the clinical workflow in radiation oncology. RO-LMM is adept at performing a variety of clinical tasks: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes. Additionally, it is capable of (2) proposing appropriate treatment plans from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed treatment plans. This multifaceted functionality demonstrates a significant alignment with the expertise of clinical professionals. To the best of our knowledge, our prototype for assisting in the radiation oncology workflow is the first work. As for an initial trial, we opted for a phased approach, starting with high prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging, i.e., breast cancer.
In the course of the aforementioned sequential tasks, the inevitable accumulation of errors in generations arises on both treatment plan suggestion and target volume segmentation task. To address this issue, we firstly explore Noisy Embedding Fine-Tuning (NEFTune) [19] which involves injecting uniform noise into embeddings during the training for our targeted tasks. To further enhance the model’s applicability, we introduce a novel consistency regularization method, bolstered by our pioneering Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs. Expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG). These advancements collectively contribute to a marked improvement in the model’s generalization capabilities in both internal and external validation. Our contributions can be summarized as:
  • We propose a comprehensive framework, denoted as RO-LMM, wherein LMM assists the entire workflow of radiation oncology for breast cancer treatment.
  • We explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG.
  • Through experiments on both internal and external datasets on real clinical breast cancer data, we demonstrated that RO-LMM outperforms baseline methods.
2 Related Works
2.1 Instruction Fine-tuning
Instruction fine-tuning has been emerged as a pivotal technique for augmenting model responsiveness in the field of natural language processing (NLP). The simplicity and effectiveness of this approach have been introduced in numerous works, with a notable scale-up facilitated by advancements of large language models (LLMs) such like GPT-3 [2], ChatGPT [36], and GPT-4 [37]. Notably, a pioneering contribution, Self-Instruct [47], involves fine-tuning of foundation models using instruction-output pairs generated from InstructGPT [38]. This methodology, along with similar approaches, has led to diverse language model variants, including Alpaca [42], Vicuna [4], Dolly [8], and a series of LLaMA [43, 44], by exhibiting promising performance across a wide range of tasks. In the medical domain, Chat-Doctor [52], Med-Alpaca [14], PMC-LLaMA [49], and Asclepius [24] have been fine-tuned for clinical question-answering (QA) task. Although these models demonstrate robust performance for diverse QA benchmarks, there is a notable gap in their applicability for the practical clinical workflow in a specific domain, such as radiation oncology, which serves as our targeted focus.
2.2 Stabilizing LLM Fine-tuning with Noise
To enhance the robustness of language models against noisy input, various strategies have been explored. Approaches such as SMART [20], FreeLB [54], and R3F [1] employ adversarial training methods, introducing small Gaussian perturbations in embedding dimensions to optimize the model’s performance against noise. Aparting from the adversarial training, LSNR [17] takes a different approach by directly optimizing the smoothness of each layer of the language model. More recently, NEFTune [19] improves LLM performance during fine-tuning by simply adding random noise into the embedding vectors during training. In the context of this study, we extend NEFTune by incorporating the concept of consistency regularization. This extension aims to further improve the robustness and generalization capabilities of language models when faced with noisy input.
Figure 2: Training schematics of the multi-purpose RO-LMM for three different tasks. (a) RO-LMM-S for clinical note summarization. (b) RO-LMM-P++ for radiation treatment plan suggestion. (c) RO-LMM-SEG++ for plan-guided target volume segmentation.
2.3 Language-driven Image Segmentation
Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [26], open-vocabulary segmentation [11, 51, 29], referring segmentation [48], and reasoning segmentation [25]. Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable. For instance, LViT [28] and ConTEXTualNet [18] introduce text-driven chest X-ray radiography segmentation.
Target volume segmentation in the field of radiation oncology is more challenging, due to its intrinsic need for considerations of the clinical aspects beyond the image, such as overall cancer stage, treatment aim, pathological findings, and so on [5, 6, 15]. Recent research on clinical context-aware breast cancer radiotherapy delineation has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce [35]. Unlike previous approaches, we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
3 Methods
In this section, we provide a detailed methodology for our proposed approach, which is designed for sequential text generation tasks, including summarization and suggestions, as well as text-driven image segmentation, as illustrated in Figure 2. In Sec 3.1, we describe text-related fine-tuning approaches that yield report summarization and plan suggestion, respectively. In Sec 3.2, we introduce a novel segmentation approach that involves injecting noise and regularizing consistency to the text embedding for improving robustness of plan-guided 3D target volume segmentation.
3.1 Consistency Embedding Fine-tuning for Clinical LMM
To realize the multi-purpose LMM with expertise in clinical report summarization and radiation treatment plan suggestion, we conduct instruction fine-tuning for LLaMA2 [44]. Considering the intended objective of each task, we adopt separate strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert). We train a summary expert using collected notes. During inference, the summary expert receives raw clinical reports, just like in the training scenario. However, for the plan expert, there is a discrepancy between the training and inference scenarios. In other words, we opt to use the training set made up of collected notes instead of generated notes, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure 1. In contrast to training, our model takes the generated notes as input for inference. To enhance robustness for the noisy input, Noisy Embedding Fine-Tuning (NEFTune) [19] which inject uniform noise into embedding is also an effective naive solution to handle noisy inputs in this task. However, a crucial consideration arises from the nature of the generated notes, which may lie closer to clean inputs (collected notes) or deviate towards noisy inputs. To address this, it is essential to train the model to handle both clean and noisy inputs. To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++†$\dagger$†$\dagger$†$\dagger$+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively. More details are as follows.
Revisiting Noise Embedding Fine-Tuning. NEFTune has been recently introduced to enhance the performance of instruction fine-tuning. This approach involves injecting a random noise vector into embeddings during the training process as follows:
$\mathcal{L}_{\text{NEFTune}}(\theta)=\mathbb{E}_{{({\mathbf{x}^{\text{emb}}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})\sim D}}\mathcal{L}_{\text{ce}}(f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}),{\mathbf{y}_{\text{txt}}}),$
${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}={\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\boldsymbol{\epsilon,\quad\epsilon}\sim\mathcal{U}(-1,1)$ (1)
where ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$ is the embedding of data, $B$ denotes batch size, $L$ is token length, $C$ is embedding dimension, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is perturbed embedding, $f_{\theta}(\cdot)$ represents the model parameterized by $\theta$, ${\mathbf{y}_{\text{txt}}}$ is the label of text sample. They demonstrated that the effectiveness of incorporating a random noise vector adjusted based on token length, which yields robust results in most fine-tuning scenarios.
Consistency Embedding Fine-Tuning. To bolster the model’s capability in handling both noisy and clean inputs for treatment plan suggestion task, we incorporate a regularization loss to encourage consistency as follows:
$\mathcal{L}_{\text{CEFTune}}(\theta)$ $=$ $\mathcal{L}_{\text{NEFTune}}(\theta)+\lambda\mathcal{R}(\theta,\theta^{-}),$ (2)
$\text{where}\,\mathcal{R}(\theta,\theta^{-})$ $=$ $d(\mathcal{T}(f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})),\mathcal{T}(f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})))$
where $\theta^{-}$ represents the model with stopped gradient, and $d(\cdot,\cdot)$ is used to quantify the discrepancy between $f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ and $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$. The function $\mathcal{T}$ serves as either the identity function or performs detokenization to generate sentences. Our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs. The metric function $d$ can be any capable of measuring distance in a specific space. However, solely minimizing distance in the same embedding space might weaken the robustness effect and pose limitations on enforcing semantic similarity. In our approach, to capture textual similarity between sentences, we calculate the distance in the feature space of SentenceBERT [40]. i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [40], $\mathcal{T}$ is set to detokenization operator. We provide algorithmic descriptions of CEFT in Supplementary Material.
3.2 LMM-assisted 3D Target Volume Segmentation
For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LMM for context-aware 3D segmentation framework introduced in [35]. However, as the LMM-assisted 3D segmentation model gets the generated treatment plan as text conditions which are inevitably noisy from the previous summarization step, we introduce noise augmentation and consistency regularization modules for the target volume segmentation network, namely, Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG).
Noise Embedding Segmentation (NESEG). Firstly, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, and denote it as NESEG. The loss function for NESEG can be formulated as:
$\mathcal{L}_{\text{NESEG}}(\Theta)$ $=\mathbb{E}_{({\mathbf{x}_{\text{img}}},{\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}},{\mathbf{y}_{\text{img}}})\sim D}\mathcal{L}_{\text{ce}}(g_{\psi}({\mathbf{x}_{\text{img}}};\mathbf{{\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}}),{\mathbf{y}_{\text{img}}})$
$\text{where}\;{\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}(\phi)=f_{\theta^{*}}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi})\,$ (3)
where $\Theta$ = [$\psi,\phi$], ${\mathbf{x}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D CT scan, $B$ denotes batch size, $H$, $W$, and $S$ correspond to height, width, and slice of the CT scan, ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ is the perturbed text embedding, ${\mathbf{y}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D ground truth segmentation mask, $g_{\psi}(\cdot)$ represents the multimodal segmentation model parameterized by $\psi$, $f_{\theta^{*}}$ denotes a frozen LLM, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is the noise injected plan embeddings, and $\mathbf{z}_{\phi}$ is the learnable text prompt parametreized by $\phi$. We denote this model as RO-LMM-SEG+.
Consistency Embedding Segmentation (CESEG). We further adapt the consistency regularization module for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan, and denote it as CESEG. Apart from the original concept of CEFTune, we modified the consistency regularization module for the multimodal segmentation task, which combines CEFTune and the text prompt tuning. Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is regularized with that of clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. The loss function for CESEG is formulated as:
$\mathcal{L}_{\text{CESEG}}(\Theta)$ $=$ $\mathcal{L}_{\text{NESEG}}(\Theta)+\lambda\mathcal{R}(\phi,\phi^{-})$ (4)
$\text{where}\,\mathcal{R}(\phi,\phi^{-})$ $=$ $d({\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}(\phi),{\mathbf{y}^{\text{emb}}_{\text{txt}}}(\phi^{-}))$
where, $\phi^{-}$ represents the learnable text prompts with stopped gradient, $d(\cdot,\cdot)$ denotes the cosine similarity, and ${\mathbf{y}^{\text{emb}}_{\text{txt}}}=f_{\theta^{*}}({\mathbf{x}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi^{-}})\in\mathbb{R}^{B\times 1\times C}$ is the text embedding for the clean plan ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. The algorithm for CESEG is provided in Supplementary Material.
Figure 3: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module. These module outputs are aligned through (c) multimodal alignment module.
The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 3. We employ the 3D Residual U-Net [7] as for the backbone structure within the image module, as shown in (a). In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input. In the multimodal alignment module, as illustrated as (c), we adopt both self-attention and cross-attention mechanisms in an interactive manner (image-to-text and text-to-image) for each $l$-th skip-connection layer, by following the concept of promptable segmentation borrowed from Segment Anything Model (SAM) [23]). For realizing multimodal alignment within the 3D segmentation framework, we adapt the idea of light-weight text prompt tuning to transfer the linguistic capability of LLM, by following the concepts introduced in [35, 22, 53]. Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module. Further details of each module are deferred to Supplementary Material.
4 Experiments
4.1 Dataset
To train our model, we collected internal data cohort composed of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center. For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 599 patients with their 3D CT scans and corresponding ground truth masks. To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital. The detail of dataset is described in Table 1 and Supplementary Material. We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals. This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
Table 1: Details of data for training our multi-purpose RO-LMM. “US" denotes Ultrasound. “Path" indicates Pathology.
Task Input Output Internal Dataset External Dataset
Train Validation Validation
Clinical Report Summarization MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
Treatment Plan Suggestion Clinical Note Treatment Plan
Plan-guided Target Segmentation Treatment Plan &\&& CT Scan Segmentation 593 79 81
4.2 Implementation details
[…]
4.3 Evaluation Metric
The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [30]. However, acknowledging potential limitations in domain specificity for the treatment plan suggestion task, we conduct expertise-based evaluations involving two clinical experts and GPT-4.0. A board-certified expert curates a total of 5 scoring rubrics for evaluation, with detailed information available in Table S9. Using the same scoring rubrics, two clinical experts assess the performance on the plan suggestion task against ground truth labels, assigning a score of 0 or 1 for each rubric separately. Inspired by the recent evaluation methodologies such as GPT-Score [13] and G-Eval [32], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning.
To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [9] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
To further validate effectiveness of our experiment, we provide the confidence intervals (CI) on each metric by utilizing the non-parametric bootstrap method. We calculate CI by conducting 1,000 times of random sampling procedure with replacement from the experimental results, each time equal to the total size of the sample dataset. Subsequently, we estimate the 95th percentile CI from the relative frequency distribution of each experimental trial.
4.4 Baseline Models
For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [14], ChatDoctor [52], Asclepius [24], and PMC-LLAMA [49]. Additionally, we include a comparison with ChatGPT [36] utilizing few-shot in-context learning for both summary and treatment plan suggestion task.
For the segmentation task, we utilize 3D Residual U-Net [7] and ConTEXTualNET [18] as our baseline methods. We further utilize an additional baseline as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
Table 2: The quantitative results for clinical note summarization using different methods. Vanilla denotes the instruction fine tuning. CI denotes confidence interval.
Model Method Internal Validation (N=60) External Validation (N=81)
Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$ (CI) Rouge L $\uparrow$ (CI) Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$(CI) Rouge L $\uparrow$ (CI)
Clinical Report Summarization
LLaMA2 [44] - 0.625 (0.603-0.646) 0.439 (0.423-0.454) 0.625 (0.603-0.646) 0.667 (0.649-0.684) 0.468 (0.456-0.479) 0.667 (0.649-0.684)
MedAlpaca [14] - 0.529 (0.484-0.578) 0.367 (0.334-0.399) 0.529 (0.484-0.578) 0.414 (0.362-0.462) 0.329 (0.283-0.373) 0.414 (0.362-0.462)
ChatDoctor [52] - 0.617 (0.572-0.657) 0.411 (0.381-0.439) 0.617 (0.572-0.657) 0.549 (0.508-0.588) 0.425 (0.392-0.459) 0.549 (0.508-0.588)
Asclepius [24] - 0.662 (0.640-0.682) 0.447 (0.436-0.458) 0.662 (0.640-0.682) 0.590 (0.571-0.609) 0.412 (0.401-0.423) 0.590 (0.571-0.609)
PMC-LLaMA [49] - 0.398 (0.341-0.451) 0.239 (0.200-0.276) 0.398 (0.341-0.451) 0.382 (0.327-0.434) 0.235 (0.197-0.275) 0.382 (0.327-0.434)
ChatGPT - 0.672 (0.641-0.703) 0.467 (0.447-0.486) 0.672 (0.641-0.703) 0.651 (0.615-0.683) 0.456 (0.434-0.476) 0.651 (0.615-0.683)
RO-LMM-S Vanila 0.798 (0.777-0.818) 0.616 (0.598-0.635) 0.798 (0.777-0.818) 0.788 (0.769-0.806) 0.607 (0.586-0.628) 0.788 (0.769-0.806)
Therapy Plan Suggestion
LLaMA2 [44] - 0.077 (0.071-0.083) 0.038 (0.034-0.042) 0.077 (0.071-0.083) 0.069 (0.065-0.073) 0.029 (0.027-0.031) 0.069 (0.065-0.073)
MedAlpaca [14] - 0.139 (0.126-0.152) 0.053 (0.049-0.058) 0.139 (0.126-0.152) 0.137 (0.123-0.152) 0.048 (0.044-0.053) 0.137 (0.123-0.152)
ChatDoctor [52] - 0.214 (0.194-0.234) 0.082 (0.072-0.091) 0.214 (0.194-0.234) 0.191 (0.178-0.205) 0.065 (0.060-0.071) 0.191 (0.178-0.205)
Asclepius [24] - 0.176 (0.162-0.190) 0.075 (0.067-0.083) 0.176 (0.162-0.190) 0.142 (0.134-0.149) 0.055 (0.051-0.059) 0.142 (0.134-0.149)
PMC-LLaMA [49] - 0.118 (0.101-0.137) 0.034 (0.029-0.040) 0.118 (0.101-0.137) 0.165 (0.136-0.195) 0.050 (0.040-0.060) 0.165 (0.136-0.195)
ChatGPT - 0.453 (0.412-0.488) 0.215 (0.187-0.246) 0.453 (0.412-0.488) 0.390 (0.357-0.423) 0.193 (0.171-0.215) 0.390 (0.357-0.423)
RO-LMM-P Vanila 0.651 (0.617-0.689) 0.486 (0.437-0.539) 0.651 (0.617-0.689) 0.591 (0.574-0.609) 0.426 (0.403-0.452) 0.591 (0.574-0.609)
RO-LMM-P+ NEFT 0.639 (0.607-0.673) 0.467 (0.419-0.516) 0.639 (0.607-0.673) 0.588 (0.573-0.606) 0.419 (0.399-0.441) 0.588 (0.573-0.606)
RO-LMM-P++ CEFT 0.668 (0.632-0.705) 0.512 (0.460-0.564) 0.668 (0.632-0.705) 0.615 (0.592-0.638) 0.459 (0.429-0.489) 0.615 (0.592-0.638)
4.5 Results
Clinical Report Summarization. We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table 2. Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals. Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning. We also perform expert analysis on the summarized reports, and the results and experimental details are deferred to Supplementary Material.
The qualitative assessment in Table 4 further highlights the effectiveness of our proposed model, providing well-organized content and consistent formatting compared to the ground truth label. While there were minor errors observed in terms of the specific distance measurement from the nipple (red), which appears to stem from the process of integrating ultrasound and MRI images, our model accurately summarizes information in a format consistent with the provided labels. In contrast, Chat-GPT exhibits issues such as omitting essential information and generating hallucinated content (red).
Treatment Plan Suggestion. As indicated in Table 2, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes. Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines. Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach. However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
Furthermore, for the treatment plan suggestion task, we comprehensively evaluate the generated plans using our proposed expertise-based rubrics as shown in Table S9. Table 3 presents the clinical expert analysis results, where R##\## represents each rubric and C##\## represents each clinical expert. For both internal and external validation, our RO-LMM variants significantly outperform the results of GPT4.0 with few-shot in-context learning due to the specific knowledge of the domain. We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals. We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$). The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
For a more detailed understanding, examples of the treatment plan suggestion task are provided in Table 4. The proposed RO-LMM-P++ (Ours) consistently yields correct answers compared to the labeled ground truth. These results demonstrate the effectiveness of our proposed CEFTune for this specific task. While Chat-GPT has shown success in providing some level of information regarding where to treat, it falls short when it comes to suggesting the specific treatment regions and dose scheme (red).
Table 3: Clinical expert analysis for treatment plan suggestion. R# represents each rubric and C# represents each clinical expert.
Internal Validation (N=25)
GPT4.0 RO-LMM-P RO-LMM-P+ (NEFT) RO-LMM-P++ (CEFT)
Experts R1 / R2 / R3 / R4 / R5 / Total R1 / R2 / R3 / R4 / R5 / Total R1 / R2 / R3 / R4 / R5 / Total R1 / R2 / R3 / R4 / R5 / Total $r$
C1 8.8 / 8.4 / 4.4 / 7.2 / 2.0 / 30.8 5.6 / 10.0 / 8.8 / 9.2 / 9.2 / 42.8 5.6 / 9.6 / 8.8 / 9.2 / 9.2 / 42.4 4.8 / 10.0 / 8.8 / 9.2 / 10.0 / 42.8 0.7
C2 9.6 / 9.6 / 5.6 / 4.4 / 3.8 / 32.7 5.6 / 10.0 / 8.4 / 8.8 / 9.2/ 42.0 5.6 / 9.6 / 8.0 / 8.4/ 10.0 / 41.6 4.4 / 10.0 / 8.8 / 9.2 / 9.6 / 42.0 0.4
GPT 10.0 / 7.6 / 4.8 / 5.6 / 6.8 / 29.2 8.2 / 8.0 / 7.2 / 8.8 / 9.6 / 37.2 9.1 / 7.6 / 6.4 / 7.6 / 9.2 / 34.8 9.1 / 8.4 / 8.0 / 8.8 / 10.0 / 39.2
External Validation (N=25) $r$
C1 8.5 / 8.8 / 7.3 / 6.5 / 1.2 / 32.3 8.1 / 8.8 / 6.9 / 6.9 / 9.2 / 40.0 7.7 / 9.6 / 7.7 / 7.3 / 9.6 / 41.9 8.5 / 9.6 / 8.8 / 9.6 / 9.6 / 46.2 0.9
C2 8.8 / 8.5 / 6.2 / 3.4 / 7.3 / 34.2 8.5 / 9.6 / 7.7 / 7.3 / 9.2 / 42.3 8.1 / 9.6 / 8.1 / 8.8 / 9.2 / 43.8 8.5 / 9.6 / 8.8 / 8.8 / 9.6 / 45.8 0.4
GPT 8.8 / 6.9 / 6.2 / 5.8 / 4.2 / 31.9 8.1 / 7.3 / 6.2 / 1.5 / 8.8 / 31.9 8.5 / 8.1 / 5.0 / 1.2 / 9.2 / 31.9 8.5 / 8.1 / 6.5 / 3.8 / 9.2 / 36.2
Table 4: Example case for the clinical report summarization task and plan suggestion task compard to baseline methods. The red scripts indicate incorrect or redundant information, and inconsistent formatting compared to the label, evaluated by a board certified clinical expert.
Clinical Report Summarization Example
User
Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
##\## Input: [MRI reports ]…[Pathology reports] …
Ground Truth
[…]
Ours
[…]
Chat-GPT
[…]
Treatment Plan Suggestion Example
User
Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
##\## Input: [Clinical notes]…
Ground Truth
RT to Lt. whole breast + IMN + SCL + AXL 40.05 Gy / 15 fx $\rightarrow$ 10 Gy boost (or 48 Gy / 15 fx with IMRT).
Ours
RT to Lt. whole breast, IMN, SCL, AXL 40.05 Gy / 15 fx + 10 Gy boost (IMRT: 48 Gy / 15 fx).
Med-Alpaca
..Breast, LLO 4:00 3cm FN, IDC, icT2N1M0, stage IIB $\rightarrow$ ypT1cN1M0, stage IIA…
ChatDoctor
…The T-stage, N-stage, ER/PR/HER2/Ki67, and Lum A status of the tumor, as well as the patient’s chemotherapy and surgical history, can all be used to determine appropriate radiation …
Chat-GPT
RT to Left breast, chest wall, L axilla Conventional fractionation 40.05 Gy / 15 fx.
Table 5: Comparison of 3D target volume segmentation performance.
Model LMM Tuning Internal Validation (N=79) External Validation (N=81)
Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
3D Residual U-Net [7] - 0.804 (0.777-0.831) 0.688 (0.653-0.720) 9.018 (7.077-11.170) 0.690 (0.652-0.725) 0.550 (0.511-0.588) 25.057 (22.844-27.070)
$\text{ConTEXTualNET}^{\dagger}$ [18] - 0.812 (0.783-0.838) 0.698 (0.661-0.731) 7.392 (5.793-9.083) 0.696 (0.663-0.726) 0.551 (0.517-0.583) 21.253 (20.002-22.474)
LLaMA2-SEG - 0.817 (0.790-0.841) 0.705 (0.670-0.734) 9.101 (7.189-11.129) 0.692 (0.666-0.714) 0.539 (0.514-0.560) 24.087 (22.450-25.789)
RO-LMM-SEG Vanilla 0.824 (0.799-0.846) 0.713 (0.680-0.742) 5.528 (3.978-7.295) 0.778 (0.743-0.806) 0.655 (0.621-0.686) 17.459 (15.254-19.678)
RO-LMM-SEG+ NESEG 0.829 (0.805-0.851) 0.719 (0.690-0.749) 4.636 (3.351-6.073) 0.781 (0.741-0.815) 0.664 (0.626-0.699) 29.410 (27.230-31.386)
RO-LMM-SEG++ CESEG 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
${}^{\dagger}$modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
Table 6: Comparison of 3D target segmentation performance between unimodal and multimodal models for both overall and specific patient types.
Types Method Internal Validation External Validation
#Patients Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) #Patients Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
All Unimodal 79 0.804 (0.777-0.831) 0.688 (0.653-0.720) 9.018 (7.077-11.170) 81 0.690 (0.652-0.725) 0.550 (0.511-0.588) 25.057 (22.844-27.070)
Multimodal 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
Gain $\uparrow$ +0.036 +0.046 -4.627 +0.076 +0.092 -9.821
Typical Unimodal 55 0.842 (0.821-0.861) 0.734 (0.705-0.761) 7.374 (5.447-9.687) 73 0.723 (0.694-0.749) 0.579 (0.546-0.610) 25.082 (22.783-27.380)
Multimodal 0.877 (0.866-0.888) 0.784 (0.766-0.802) 3.200 (2.361-4.303) 0.782 (0.742-0.813) 0.659 (0.621-0.690) 14.716 (12.407-17.204)
Gain $\uparrow$ +0.035 +0.049 -4.173 +0.059 +0.080 -10.366
Atypical Unimodal 24 0.720 (0.654-0.784) 0.584 (0.513-0.657) 12.519 (8.553-16.584) 8 0.370 (0.178-0.594) 0.259 (0.106-0.449) 27.138 (22.417-30.807)
Multimodal 0.760 (0.701-0.807) 0.628 (0.562-0.682) 6.948 (3.820-11.014) 0.585 (0.346-0.753) 0.458 (0.269-0.610) 22.410 (16.377-28.314)
Gain $\uparrow$ +0.040 +0.044 -5.571 +0.215 +0.199 -4.728
3D Target Volume Segmentation. As reported in Table 5, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting. This can be also observed in qualitative results. As shown in Figure 4(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks. Moreover, as shown in Figure 4(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows. RO-LMM-SEG without any proposed module shows reasonable performance, yet sometimes undersegments the treatment target volume, as shown in Figure 4(b). In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated.
We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type). The comparative results are presented in the Table 6 separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG. Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively. When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%. This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
Figure 4: Qualitative comparison on 3D target volume segmentation task between unimodal and multimodal approaches. Red arrows indicate incorrectly segmented regions.
5 Discussion
5.1 Analysis on Strategy of Separate Expert
To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing a unified model shown in Table 8. We observe that the unified model’s performance is significantly inferior to our distinct training strategy. This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task. These results suggest that the knowledge required for each summarization and suggestion task necessitates task-aware training strategies.
Table 7: Ablation study on adopting separate expertise for each textual task against unified strategy.
Table 8: Ablation study on CESEG with input text variation for target segmentation performance.
Training Strategy Internal (N=60) External (N=81)
R-1 / R-2 / R-L R-1 / R-2 / R-L
Clinical Note Summary
Unified strategy 0.516 / 0.357 / 0.515 0.327 / 0.215 / 0.326
Distinct strategy 0.823 / 0.635 /0.822 0.805/ 0.634 / 0.804
Therapy Plan Suggestion
Unified strategy 0.510 / 0.265 / 0.502 0.409 / 0.197 / 0.406
Distinct strategy 0.652 / 0.491 / 0.649 0.606 / 0.450 / 0.606
CESEG Input Text Target Segmentation
Dice $\uparrow$ IoU $\uparrow$
Ground Truth $0.860_{\pm 0.073}$ $0.761_{\pm 0.103}$
Generated $0.825_{\pm 0.101}$ $0.714_{\pm 0.134}$
Difference $\downarrow$ $0.035$ $0.047$
Ground Truth $0.847_{\pm 0.088}$ $0.743_{\pm 0.111}$
Generated ${0.840_{\pm 0.094}}$ ${0.733_{\pm 0.120}}$
Difference $\downarrow$ 0.007 0.010
Table 8: Ablation study on CESEG with input text variation for target segmentation performance.
5.2 Analysis on Consistency Regularization
[…]
5.3 Limitation
While our study provides valuable insights on the application of radiation oncology, there are several limitations to consider. Firstly, the dataset used in this study is currently confined to patients with their initial diagnoses, primarily focusing on breast cancer. Expanding the scope of our research to include diverse patient scenarios and other cancer types is essential for enhancing the applicability of our model. Secondly, achieving detailed target volume segmentation for Simultaneous Integrated Boost (SIB) necessitates information beyond treatment plan-guided segmentation. This additional information may be obtained from alternative imaging modalities, such as MRI or PET-CT, along with ground truth masks for organ-at-risks (OARs) and gross tumor volume (GTV). Consequently, SIB is excluded from this study. Future research will focus on developing a model for SIB by incorporating a more extensive array of information.
6 Conclusion
In this work, we introduce RO-LMM, a multi-purpose foundation model tailored for radiation oncology. Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation, mirroring real-world clinical workflows. Another key contribution of this work is the introduction of consistency technique into both text and segmentation task. Results from multi-center cohort datasets confirm RO-LMM’s promising performance and noteworthy generalization capabilities across diverse tasks. These findings mark a significant stride toward developing a versatile AI model, hinting at the potential for a multi-purpose medical AI model in radiation oncology.
Ethical Statement All the patient information and any potential personally identifiable information from used datasets are de-identifiied, and clinical data usage is approved by the Institutional Review Board (IRB).
Appendix 0.A Supplementary Section
In this supplemental documents, we present:
  • An algorithmic descriptions of our proposed Consistency Embedding Fine Tuning (CEFT) and Consistency Embedding Segmentation (CESEG) in Section 0.B.
  • An explanation of the network architecture for the segmentation model in Section 0.C.
  • Results of ablation studies for both text-related tasks and target volume segmentation tasks in Section 0.D and Section 0.C.
  • Detailed information about the dataset, including pre-processing and acquisition procedures, in Section 0.F.
  • Examples of the entire workflow in Section 0.I.
  • Additional comparison results for all tasks in Section 0.G.
  • Detailed descriptions of the evaluation for treatment plan suggestion, including prompts for evaluation, in Section 0.J.
  • Detailed descriptions of the expert-analysis for clinical report summarization, including prompts for evaluation, in Section 0.H.
Appendix 0.B Algorithm of CEFTune and CESEG
Algorithm 1 and Algorithm 2 details the training procedure with CEFTune and CESEG, respectively.
Input :  $\{X^{\text{txt}}_{i},Y^{\text{txt}}_{i}\}_{i=1}^{N}$ $\sim D$ tokenized dataset, large language model $f_{\theta}(\cdot)$, detokenization operator $\mathcal{T(\cdot)}$, the metric which calculate the cosine similiary $d(\cdot,\cdot)$, the hyper-parameter for the noise intensity and loss, $\alpha$, and $\lambda$, repectively. The cross entropy loss $\mathcal{L}_{\text{ce}}$,
the embedding layer $\text{emb}(\cdot)$, $\theta^{-}$leftarrow stopgrad $(\theta)$;
1 repeat
2       $\text{Sample}\;({\mathbf{x}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})$ $\sim D$ ;
3       ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\leftarrow\text{emb}({\mathbf{x}_{\text{txt}}})$ ;
4       ${\mathbf{\hat{y}}_{\text{txt}}}\leftarrow f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$ ;
5       $\epsilon\sim\mathcal{U}(-1,1)$;
6       ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}\leftarrow{\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\epsilon$ ;
7       ${\mathbf{\tilde{y}}_{\text{txt}}}\leftarrow f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ ;
8       $\mathcal{L}(\theta)\leftarrow$ $\mathcal{L}_{\text{ce}}({\mathbf{\hat{y}}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})+\lambda d(\mathcal{T}({\mathbf{\tilde{y}}_{\text{txt}}}),\mathcal{T}({\mathbf{\hat{y}}_{\text{txt}}}))$
9until Stopping criteria met max iterations;
Algorithm 1 CEFTune: Consistency Embedding Fine tuning
Input :  $\{X_{i}^{\text{img}},X_{i}^{\text{txt}},Y_{i}^{\text{img}}\}_{i=1}^{M}$ $\sim D$, the entire image segmentation model $g_{\Theta}(\cdot)$, where $\Theta$ = [$\psi,\phi$], which consist of segmentation model $g_{\psi}$, and learnable text prompt $z_{\phi}$. The frozen large language model $f_{\theta^{*}}(\cdot)$, output embedding of $f_{\theta^{*}}$ ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$, the hyper-parameter for the noise intensity and loss $\alpha$ and $\lambda$, respectively, the metric which calculate the cosine similarity $d(\cdot,\cdot)$, the cross entropy loss $\mathcal{L}_{\text{ce}}$,
the embedding layer $\text{emb}(\cdot)$, $\phi^{-}$leftarrow stopgrad $(\phi)$,
1 repeat
2       $\text{Sample}\;({\mathbf{x}_{\text{img}}},{\mathbf{x}_{\text{txt}}},{\mathbf{y}_{\text{img}}})$ $\sim D$ ;
3       ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\leftarrow\text{emb}({\mathbf{x}_{\text{txt}}})$ ;
4       ${\mathbf{y}^{\text{emb}}_{\text{txt}}}\leftarrow f_{\theta^{*}}({\mathbf{x}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi})$ ;
5       $\epsilon\sim\mathcal{U}(-1,1)$;
6       ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}\leftarrow{\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\epsilon$ ;
7       ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\leftarrow f_{\theta^{*}}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi^{-}})$ ;
8       $\mathcal{L}(\Theta)\leftarrow$ $\mathcal{L}_{\text{ce}}(g_{\psi}({\mathbf{x}_{\text{img}}};\mathbf{{\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}}),{\mathbf{y}_{\text{img}}})+\lambda d({\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}},{\mathbf{y}^{\text{emb}}_{\text{txt}}})$
9until Stopping criteria met max iterations;
Algorithm 2 CESEG: Consistency Embedding Segmentation
Appendix 0.C Architecture of RO-LMM-SEG++
The schematic of our RO-LMM-SEG++ is illustrated in Figure 3. For text prompt tuning of LLM, we introduce a total of $N$-text prompts $\{{{\mathbf{z}_{\phi}}^{n}|^{N}_{n=1}}\}$, where each ${\mathbf{z}_{\phi}}^{n}\in\mathbb{R}^{M\times C}$ is composed of $M$ vectors with the dimension $C$. These learnable vectors are randomly initialized, and then consistently prepended to each of treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$. Here, the final prompted text input $t$ can be formulated as follows:
$t=\{{\mathbf{z}_{\phi}}_{1}^{n},{\mathbf{z}_{\phi}}_{2}^{n},...,{\mathbf{z}_{\phi}}_{M}^{n},{\mathbf{x}^{\text{emb}}_{\text{txt}}}\}.$ (5)
Then, using the prompted text input $t$, the frozen LLM resulted the text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ as for the last token output. Then, to align the text embedding with the intermediate image embedding ${h}_{img}^{l}\in\mathbb{R}^{B\times H_{l}W_{l}S_{l}\times Ch_{l}}$, ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ is projected to yield ${h}_{txt}^{l}\in\mathbb{R}^{B\times 1\times Ch_{l}}$ via $l$-th layer-wise linear layer, where $H_{l}$, $W_{l}$, and $S_{l}$ correspond to height, width, and slice of the image embedding, and $Ch_{l}$ is the channel dimension of $l$-th skip-connection layer within the image module. For multimodal alignment, the projected text embeddings ${h}_{txt}^{l}$ is self-attended and crossly-attended with the image embedding ${h}_{img}^{l}$ for each layer, and each multimodal embedding ${h^{*}}^{l}_{img}$ is inputted to the decoder part of the image module to yield final prediction $\hat{y}_{\text{img}}\in\mathbb{R}^{B\times 1\times HWS}$.
Table S1: Component analysis of RO-LMM-P framework on treatment plan suggestion.
Component Condition Ours Internal Validation (N=60) External Validation (N=81)
Rouge 1 $\uparrow$
Rouge 2 $\uparrow$
Rouge L $\uparrow$
Rouge 1 $\uparrow$
Rouge 2 $\uparrow$
Rouge L $\uparrow$
Teacher Module
Noisy teacher
0.613
0.453
0.615
0.575
0.428
0.571
Clean teacher
$\checkmark$
0.652 0.491 0.649 0.606 0.450 0.604
SBERT Moudle
SBERT
0.648
0.484
0.645
0.598
0.442
0.594
PubMedBERT
$\checkmark$
0.652 0.491 0. 649 0.606 0.450 0.604
Noise Intensity ($\alpha$) 5
$\checkmark$
0.652 0.491 0.649 0.606 0.450 0.604
10
0.648
0.488
0.647
0.597
0.435
0.598
15
0.642
0.481
0.642
0.576
0.415
0.572
Noise type Uniform 5
$\checkmark$
0.652 0.491 0.649 0.606 0.450 0.604
Gaussian 5
0.625
0.449
0.629
0.588
0.425
0.586
Backbone
LLaMA-13B
0.645
0.478
0.640
0.597
0.436
0.594
+NEFT
0.658
0.500
0.653
0.601
0.438
0.597
+CEFT
$\checkmark$
0.662 0.502 0.661 0.613 0.441 0.603
Temperature
0.6
0.647
0.480
0.643
0.593
0.435
0.591
0.8
$\checkmark$
0.652 0.491 0.649 0.606 0.450 0.604
1.0
0.662
0.487
0.647
0.601
0.442
0.599
Appendix 0.D Ablation Study on Therapy Plan Suggestion
To investigate the effectiveness of our proposed methods, we conduct the ablation studies for treatment plan suggestion task, as shown in Table S1.
Consistency Regularization Loss We manipulate the objective function in Equation 3.1 to analyze the effect of consistency regularization. By default, we use $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$ as the target (referred to as “clean teacher"). For comparison, we set $f_{\theta^{-}}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ as the teacher and $f_{\theta}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$ as the student (referred to as "noisy teacher"). We observe that the clean teacher significantly outperforms the noisy teacher. By restricting the divergence of the distribution of students, the clean teacher yields enhanced performance compared to the noisy teacher.
Effect of SBERT Module We study the effect of the SBERT module in enforcing consistency between clean and noisy input. Our model, equipped with the medical domain-specific SBERT variant, PubMedBERT, outperforms models trained on general language. This demonstrates the effectiveness of incorporating domain-specific knowledge.
Analysis of Noise type and Intensity ($\alpha$) Since our proposed method involves injecting noise into embeddings, we conduct a study to analyze the effect of noise type and intensity under various conditions. First, regarding the noise type, we found that our method performed best when equipped with uniform noise compared to Gaussian noise. Consequently, we set the default configuration for noise type to uniform noise. Additionally, we analyze the effect of noise intensity by varying the parameter $\alpha$ from 5 to 15. Smaller values of $\alpha$ resulted in better performance on the internal dataset, while larger values improved performance on the external dataset. However, excessively large values of $\alpha$ led to a significant drop in performance on the external dataset. Based on empirical results, we select $\alpha$ set to 5 as the baseline.
Analysis of Model Backbone We conduct experiments using the larger 13B LLaMA-2 backbone version and observ a similar performance trend compared to the 7B model. Despite the larger backbone size, our proposed CEFT method outperform the other baselines, demonstrating scalability. Taking into account both cost-effectiveness and performance, we adopt the 7B model as the default configuration.
Analysis of Temperature To investigate the impact of stochasticity, we conducted experiments varying the temperature during treatment plan suggestion generation. Our findings indicate that the results produced by our proposed method are insensitive and robust to changes in temperature. Based on empirical analysis, we set the default temperature to 0.8.
Appendix 0.E Ablation Study on Target Segmentation
To optimize treatment target segmentation performance, we conduct the ablation studies, as shown in Table S2.
Analysis of LMM Tuning Methods We compare the different LMM tuning methods to find optimal performance, where NESEG nor CESEG module is not activated. We analyze various methods including no tuning, low-rank adaptation (LoRA) fine-tuning [16] and prompt tuning methods with a single or multiple text prompts. From the experimental results, the text prompt tuning shows the most reliable performance across the external validation setting. Additionally, the use of single text prompt shows a remarkable improvement in performance compared to utilizing multiple text prompts.
Analysis of Noise Intensity ($\alpha$) We analyze the effect of noise intensity by varying $\alpha$ from 5 to 20, where CESEG module is not activated. We empirically find that $\alpha$ as 10 yields the best performance. We further train the hyper-parameter $\alpha$ as a learnable parameter, however, yielding inferior performance compared to employing the fixed noise.
Table S2: Component analysis of RO-LMM-SEG framework on target volume segmentation performance.
Component Condition Ours Internal Validation (N=79) External Validation (N=81)
Dice $\uparrow$
IoU $\uparrow$
HD-95 $\downarrow$
Dice $\uparrow$
IoU $\uparrow$
HD-95 $\downarrow$
LMM Tuning Method
None
$0.822$
$0.712$
$12.368$
$0.752$
$0.622$
$24.849$
LoRA [16]
$0.810$
$0.699$
$\bf{4.989}$
$0.759$
$0.637$
$\bf{16.692}$
Text Prompts (p=1)
$\checkmark$
$\bf{0.825}$
$\bf{0.714}$
$5.392$
$\bf{0.775}$
$\bf{0.652}$
$17.623$
Text Prompts (p=2)
$0.822$
$0.711$
$7.030$
$0.757$
$0.627$
$24.64$
Noise Intensity ($\alpha$)
5
$0.821$
$0.707$
$5.917$
$0.738$
$0.601$
$\bf{22.320}$
10
$\checkmark$
$\bf{0.830}$
$\bf{0.720}$
$\bf{4.383}$
$\bf{{0.771}}$
$\bf{{0.653}}$
$28.835$
20
$0.826$
$0.716$
$5.575$
$0.732$
$0.593$
$26.090$
Learnable
$0.828$
$0.715$
$5.024$
$0.757$
$0.633$
$23.694$
Appendix 0.F Details of Training Dataset
For data preparation, we collected internal training & validation data from patients treated at Yonsei Cancer Center between January 2009 and December 2022. For external validation, we further collected data from patients from Yongin Severance Hospital between January 2018 and December 2022. We confirmed that the external test set was not overlapped with the training dataset.
For pre-processing image data for training RO-LMM-SEG++, all the 3D chest CT scans and target volumes were re-sampled with identical voxel spacing of 1.0 $\times$ 1.0 $\times$ 3.0 $\text{mm}^{3}$. The intensity values of the CT scans were truncated between -1,000 and 1,000 of Hounsfield unit (HU), and linearly normalized between 0 and 1.0. When training, a 3D patch with a spatial dimension of 384 $\times$ 384 $\times$ 128 pixels was randomly cropped. When inference, the entire 3D CT scans was tested using sliding windows with the identical spatial dimension of the 3D patch to that used for training.
Appendix 0.G Additional Qualitative Results
In this section, we provide the additional qualitative results on all tasks as shown in Table S4, Table S5 and Figure S1.
In the context of report summarization task involving cases with multiple (two or more) tumor masses, as shown in in Table S4, our model adeptly summarize the existence of two tumors in the first line of the summary, despite minor inaccuracies in pinpointing the exact locations of the tumors. In contrast, ChatGPT incorrectly summarize these cases as having only one tumor mass in the first line. Additionally, ChatGPT exhibit instances of hallucination, such as inappropriately mentioning tumor sizes in the first line, erroneously including post-therapy (yp) stages for patients who did not receive chemotherapy, and excessively condensing key information such as the presence of the lymph node metastasis, making it challenging to discern. Overall, unlike our model which aligns closely with the ground truth in its summaries, ChatGPT demonstrates significant issues, including the generation of erroneous hallucination and the failure to comprehensively encapsulate necessary details.
In the evaluation of treatment plan suggestion capabilities in Table S5, only our model and ChatGPT are able to offer somewhat meaningful suggestions related to radiotherapy treatment, while other models fail. ChatGPT, drawing upon its built-in knowledge base, seems to propose plans that are partially relevant. However, it exhibitd a fundamental misunderstanding of dose schemes, frequently misinterpreting hypofractionation as conventional fractionation. A significant limitation in ChatGPT’s approach is its inability to incorporate all regional lymph node areas into the target volume, often disproportionately focusing on specific areas, such as the axillary lymph nodes. Additionally, ChatGPT struggles to distinguish between the treatment of the breast and the chest wall, leading to flawed treatment plan suggestions. These limitations seen to arise from ChatGPT’s lack of a holistic understanding of radiation oncology, unlike our domain-specific model.
Figure S1 provides qualitative comparisons of the target volume segmentation performance for breast cancer patients, where the unimodal model fails to segment target volume with incorrect laterality as indicated as red arrows. On the other hand, our proposed multimodal RO-LMM-SEG++ with CESEG shows the most promising qualitative segmentation performance, aligning consistently with the ground truth masks.
Figure S1: 3D target volume segmentation cases compared to baseline methods. Red arrows indicate incorrectly segmented regions.
Appendix 0.H Expert Analysis on Clinical Report Summarization
Similar to treatment plan suggestions, we evaluate the generated summaries using our proposed expertise-based rubrics, as illustrated in Tab. S9, and compare them with other baselines, including ChatGPT and LLaMa-2. Tab. S3 presents the results of clinical expert analysis, where R##\## represents each rubric and C##\## represents each clinical expert. For both internal and external validation, our ROB-LMM-S significantly outperforms all other baselines due to its specific knowledge in this domain. We further evaluate inter-clinician correlation using Pearson correlation ($r$), with results showing strong positive correlations ($>0.85$ and $>0.95$, for each internal and external validation), affirming the reliability of our proposed evaluation rubrics and the clinical effectiveness of our ROB-LMM-S. Therefore, our ROB-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
Table S3: Clinical expert analysis for Clinical report summarization. R# represents each rubric and C# represents each clinical expert.
Internal Validation (N=10)
GPT LLaMa2-7B RO-LMM-S
Experts R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total $r$
C1 3.0 / 0.0 / 4.0 / 7.0 / 13.3 2.0 / 0.0 / 0.0 / 7.0 / 9.0 15.0 / 10.0 / 10.0 / 10.0 / 45.0 0.884
C2 3.0 / 0.0 / 5.0 / 7.0 / 14.4 1.0 / 0.0 / 0.0 / 9.0 / 10.0 15.0 / 10.0 / 9.0 / 10.0 / 44.0
External Validation (N=10) $r$
C1 5.0 / 1.0 / 9.0 / 7.0 / 22.0 7.0 / 0.0 / 0.0 / 7.0 / 14.0 10.0 / 9.0 / 8.0 / 8.0 / 35.0 0.977
C2 6.0 / 3.0 / 9.0 / 7.7 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
Table S4: Example case for the clinical report summarization task. The red scripts indicate the incorrect or redundant information, and inconsistent format compared to label.
Clinical Report Summarization Example
User
Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
####\#\## # Input: [MRI report][Pathology report]
Ground truth
##\## 1. Breast, RLO 08:00 4cm FN, RLM 04:00 2cm FN, IDC, two, pT1cN1aM0, stage IIA
- Histologic type: Invasive ductal carcinoma, NOS
- Tumor margin: Infiltrative
- Maximum diameter of invasive tumor: 1.5cm and 0.4cm
- Maximum diameter of invasive and in situ carcinoma: 3.0cm
- Nuclear grade: 2 - Histologic grade: II (3-2-1)
- LVI (-) PNI (-) - Intraductal carcinoma component
[…]
- Type: Cribriform - Nuclear grade: Low
- Necrosis: Absent - Location: Periphery
- Calcification: Present (Malignant abd benign)
- Associated benign lesion: Intraductal papilloma, sclerosing adenosis and columnar cell lesion
- RM (-) - Overlying skin (superficial margin): Free of ca (safety margin: 1.4cm)
- Underlying fascia (deep margin): Free of ca (safety margin: 2.3cm)
Lymph nodes, total (2/13): PNE (-, Maximal diameter: 3mm)
- Axillary (0/10): Free of ca
- Sentinel lymph node1 (1/1), axillary lymph node1 (1/1), axillary lymph node2 (0/1) and axillary lymph node3 (0/0); total (2/3): PNE (-, Maximal diameter: 3mm)
[…]
Oncotype Dx: RS 10 CTx omission
Ours
##\## 1. Breast, RLO 09:00 4cm FN RMC 03:00 3cm FN, IDC, Multifocal, pT1cN1aM0, stage IIA
- Histologic type: Invasive ductal carcinoma, NOS
- Tumor margin: Infiltrative
- Maximum diameter of invasive tumor: 1.5cm and 0.4cm
- Maximum diameter of invasive and in situ carcinoma: 3.0cm
- Nuclear grade: 2 - Histologic grade: II (3-2-1)
- Lymphovascular invasion: Absent - Perineural invasion: Absent - Intraductal carcinoma component
[…]
- Type: Cribriform - Nuclear grade: Low
- Necrosis: Absent - Location: Periphery
- Calcification: Present (Malignant abd benign)
- Associated benign lesion: Intraductal papilloma, sclerosing adenosis and columnar cell lesion
- Resection margins - Overlying skin (superficial margin): Free of carcinoma (safety margin: 1.4cm)
- Underlying fascia (deep margin): Free of carcinoma (safety margin: 2.3cm)
Lymph nodes, total (2/13): PNE (-), max 3mm
- Axillary (0/10): Free of carcinoma
- Sentinel lymph node1 (1/1), axillary lymph node1 (1/1), axillary lymph node2 (0/1) and axillary lymph node3 (0/0); total (2/3): Metastatic carcinoma in 2 out of 3 lymph nodes without perinodal soft tissue extension (Maximal diameter: 3mm)
[…]
Oncotype DX RS: 20 Adj. CTx omission
Chat-GPT
##\##1. Breast, RLO 08:00, 15mm, IDC, pT1cN1(f)M0, Stage IIA ypT2 0cN0M0
- two (1.5cm and 0.4cm), H2N2
- two of 13 +
- LVI (-), DLI (-), HD (-)
- Extensive IC: (-) - Calcification (+) - RM (-)
- Overlying skin: Free of ca (safety margin: 1.4cm)
- Underlying fascia: Free of ca (safety margin: 2.3cm)
[…]
Table S5: Example case for the plan suggestion task compared to baseline methods. The red scripts indicate the incorrect answers.
Example 1
User
Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
##\## Input: [Clinical notes]
Ground truth
RT to WB and tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx.
Ours
RT to whole breast + tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx.
Med-Alpaca
…Breast, RUC, 12:00, 5cm FN, IDC, pT1cN0M0, stage IA…
ChatDoctor
… Depending on the patient’s individual circumstances, the radiation oncologist may recommend a course of external beam radiation therapy, brachytherapy, or a combination of the two …
Chat-GPT
RT to right breast Conventional fractionation Total dose: 50 Gy in 25 fractions.
Example 2
User
Plan a radiotherapy treatment based on the insights from the radiation oncologist’s clinical note.
##\##Input: [Clinical notes]
Ground truth
RT to Rt. chest wall + AXL Lv 3, IMN, SCL 40.05 Gy / 15 fx.
Ours
RT to Rt. chest wall, AXL Lv 3, IMN, SCL 40.05 Gy / 15 fx.
Med-Alpaca
.. Status post core biopsy of the right breast, performed …
ChatDoctor
… Deliver a dose of 40 Gy in 10 fractions over 4 weeks to the primary tumor, with a margin of 1.1 cm for the overlying skin and 1.2 cm for the underlying fascia. …
Chat-GPT
RT to chest wall, IMN, and AXL Conventional fractionation 40.05 Gy / 15 fx.
Example 3
User
Considering the clinical insights from the radiation oncologist’s note, set a treatment plan in terms of radiotherapy.
##\## Input: [Clinical notes]
Ground truth
RT to Lt. WB and tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy in daily fx).
Ours
RT to Lt. WB and tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy daily fx)
Med-Alpaca
.. Lymph nodes, total (0/4): Free of carcinoma - Sentinel lymph node1 (0/1), sentinel lymph node2 (0/1), sentinel lymph node3 (0/1), and axillary lymph node1 (0/1); total (0/4): Free of carcinoma …
ChatDoctor
… Radiotherapy will be administered to the breast, left, 2 o’clock, 4 cm FN IDC, pT1bN0M0, H1N2, PNI (-), LVI (-), EIC (-), Calcification: …
Chat-GPT
RT to Left breast Conventional fractionation 40.05 Gy / 15 fx.
Example 4
User
Considering the clinical insights from the radiation oncologist’s note, set a treatment plan in terms of radiotherapy.
##\## Input: [Clinical notes]
Ground truth
RT to Lt. WB/Cavity, AXL 2, 3, Lv 4 (ESTRO), IMN 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy in daily fx).
Ours
RT to Lt. WB/Cavity, AXL 2, 3, Lv 4 (ESTRO), IMN 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy daily fx).
Med-Alpaca
[…]
ChatDoctor
… These findings suggest that the patient has an aggressive form of breast cancer and may benefit from additional chemotherapy and targeted therapy. …
Chat-GPT
RT to left chest wall and axilla Conventional fractionation 40.05 Gy / 15 fx for chest wall 40.05 Gy / 15 fx for axilla.
Appendix 0.I Examples of Entire Workflow
We present examples of the entire workflow as indicated in Table S6 and Table S7. RO-LMM can take input as a pair of MRI report and Pathology report or another pair of Ultrasound report and Pathology report. Our model seamlessly serves as an assistant for radiation oncologists by conducting report summarization, treatment plan suggestion, and 3D Target Volume Segmentation.
Table S6: Example of Entire Workflow in the case of taking MRI report and Pathology report as input.
Case 1. MRI report and Pathology report
User
Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
####\#\## # Input: [MRI report][Pathology report]
RO-LMM-S+
##\##1. Breast, ROC, 09:00, 3cm FN, IDC, icT2N1(f)M0
stage IIB (anatomic), stage IIIB (prognostic) ypT0N0M0
[ic]
T: - 39mm size mass, H3N3 - LVI (-)
N: - Multiple suspicious lymph nodes at Rt. axilla level 1, 2, 3 (FNA+)
[…]
[yp]
- No residual ca
- Fibrosis with hemosiderin laden macrophages and foam cell collection, favoring regressed tumor bed
- Total LNs (0/25) : Axillary (0/11), Level II (0/6) : SLN (0/4), axillary (non SLN) (0/4)
- Calcification (+, Benign)
- RM (-): Nipple, upper, and lower margin: Free of carcinoma (See SF21-04688)
* Lymphedema prediction
[…]
User
Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
####\#\## # Input:[Generated Clinical Note]
RO-LMM-P++
RT to Rt. whole breast + AXL Lv 3 + SCL 48 (SIB) - 40.05 Gy / 15 fx.
User
Draw the radiotherapy target volume of this patient with following planning CT scan.
####\#\## # Input:[Generated Treatment Plan][3D CT Scan] patient001.dcm
RO-LMM-SEG++
patient001__\__segmentation.dcm
Table S7: Example of Entire Workflow in the case of taking Ultrasound report and Pathology report as input.
Case 2. UltraSound report and Pathology report
User
Summarize a clinical summary note as a radiation oncologist might, using the provided Ultrasound and pathology details.
####\#\## # Input: [Ultrasound reports ][Pathology reports]
RO-LMM-S+
##\##1. Breast, LUO 02:00 2cm FN, IDC, pT1cN0M0, stage IA
- Histologic type: Invasive ductal carcinoma
- Tumor margin: Infiltrative - Nuclear grade: 2 - Histologic grade: II (3-2-1)
- Maximum diameter of invasive tumor: 1.4cm
- Maximum diameter of invasive and in situ carcinoma: 2cm
- Lymphovascular invasion: Not identified - Perineural invasion: Not identified
[…]
- Type: Solid / Cribriform - Nuclear grade: Intermediate - Necrosis: Present (Comedo)
- Location: Central and periphery - Calcification: Present (Malignant and benign)
[…]
- Resection margin: - Upper, lower, inner and outer margin: Free of ca
- Total LNs (0/11): SLN (0/6), non SLN (0/5)
ER: Positive (AS 8 = PS 5 + IS 3)
PR: Positive (AS 5 = PS 3 + IS 2)
HER2: Negative (1+)
AR: Positive (AS 7 = PS 5 + IS 2)
[…]
User
Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
####\#\## # Input:[Generated Clinical Note]
RO-LMM-P++
RT to Lt. whole breast 26 Gy / 5 fx + 10 Gy / 5 fx boost.
User
Draw the radiotherapy target volume of this patient with following planning CT scan.
####\#\## # Input:[Generated Treatment Plan][3D CT Scan] patient002.dcm
RO-LMM-SEG++
patient002__\__segmentation.dcm
Appendix 0.J Prompt for Evaluation of Treatment Plan Suggestion
In this paper, we design prompts for evaluating generated treatment plans, including a reference-guided score rubric curated by clinical expert, as shown in Table S9. Similar to previous GPT-based evaluation methods, we tailor the prompts combining the specific knowledge of clinical experts. By using this prompt, we evaluate the results of our RO-LMM and all the baseline methods using GPT-4.0.
Table S8: Used prompts with proposed expertise-based rubrics for assessing the performance of treatment plan suggestion.
Based on <Ground Truth Plan>, score 1 if <Model Generated Plan> satisfies each rubric, else score 0.
[Rubric]
R1. Laterality.
Whether orientation (Rt./Lt.) is correct
R2. Aim of Surgery.
Whether aim of RT (definitive RT = RT after breast conservation surgery or partial mastectomy. In this scenario, breast should be irradiated. / postop RT = RT after total or radical mastectomy. In this scenario, chest wall should be irradiated) is correct.
R3. Treatment Scope.
Whether detailed extent of radiation field (inclusion or exclusion of WB, CW, AXL, SCL, IMN) and detailed level of inclusion (RTOG = in this guideline, upper SCL node is included / ESTRO = in this guideline, upper SCL is not included) is correct.
R4. Dose Scheme.
Whether RT scheme (dose in Gy, and fractions in fx) is correct. It also included SIB or sequential boost dose.
R5. Hallucination.
Whether unnecessary information is added or not (If unnecessary & incorrect information was added, score 0 points, otherwise score 1 point)
[Abbreviations]
RT.
Radiotherapy
BCS.
Breast Conserving Surgery
PMRT.
Post Mastectomy Radiation Therapy
IMRT.
Intensity Modulated Radiation Therapy
PBI.
Partial Breast Irradiation
SIB.
Simultaneous Integrated Boost technique, same as tumor bed boost (sequential), fast-forward protocol
SCL.
Supraclavicular Lymph Node
IMN.
Internal Mammary Lymph node
AXL.
Axilla Lymph Node
RTOG.
Radiation Therapy Oncology Group
ESTRO.
European Socity for Radiotherapy and Oncology
CW.
Chest Wall
WB.
Whole Breast
[Ground Truth Answer Begin]
<Ground Truth Plan>
[Ground Truth Answer End]
[Model Responses Begin]
<Model Generated Plan>
[Model Responses End]
[Answer Format]
Model, Answer for R1, Answer for R2, Answer for R3, Answer for R4, Answer for R5, (scores ONLY for each Rubric
in a tabular format, without explanation):
Table S9: The proposed expertise-based rubrics for assessing the performance of clinical report summarization.
Based on rubrics, score each clinical report summarization.
[Rubric]
Rubric 1: Relevance
Assessment of whether the necessary information from the Ground Truth (GT) is included, and identification of any extraneous information
Good (2): The information from the GT is appropriately included.
Partially correct (1): Some relevant information is included, but there are omissions or irrelevant details.
Mostly irrelevant (0): The majority of the information is unnecessary or irrelevant.
Rubric 2: Formatting
Evaluation of whether the necessary information is presented in the correct format and location:
Incorrectly placed (0): Information is not presented in the appropriate format or location.
Correctly placed (1): The necessary information is accurately positioned in the expected format.
Rubric 3: Length
Assessment of whether the writing is excessively verbose:
Too verbose (0): The writing is overly lengthy and lacks conciseness.
Sufficiently summarized (1): The content is effectively summarized without unnecessary verbosity.
Rubric 4: Consistency
Identification of conflicting information between sources:
Present (0): Conflicting pieces of information are observed.
Absent (1): There are no conflicting pieces of information between sources.

### v3

@@
-LMM-Assisted Breast Cancer Treatment Target Segmentation with Consistency Embedding
+End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding
 Abstract
-Recent advancements in Artificial Intelligence (AI) have profoundly influenced medical fields, by providing tools to reduce clinical workloads.
-However, most AI models are constrained to execute unimodal tasks, in stark contrast to the comprehensive approaches utilized by medical professionals.
-To address this, here we present RO-LMM, a multi-purpose large multimodal model (LMM) tailored for the field of radiation oncology.
-This model covers series of tasks within clinical workflow, adept at clinical report summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation.
-In particular, to perform consecutive clinical tasks, we further present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the capability of handling clean inputs, and transform this concept into LMM-driven segmentation framework as Consistency Embedding Segmentation (CESEG).
-Experimental results on multi-centre cohorts demonstrate our RO-LMM’s promising performance for multiple clinical tasks with generalization capabilities.
-Figure 1: RO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology.
-The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation.
-All the patient information and any potential personally identifiable information are de-identifiied, and the clinical data usage is approved by IRB.
-**footnotetext: These authors contributed equally to this work\$$\dagger$\$$\dagger$\$footnotetext: Corresponding authors
-1 Introduction
-Recently, the emergence of a new generation of AI models, known as foundation models, marks a significant departure from previous paradigms [34].
+Recent advances in AI foundation models have significant potential for lightening the clinical workload by mimicking the comprehensive and multi-faceted approaches used by medical professionals.
+In the field of radiation oncology, the integration of multiple modalities holds great importance, so the opportunity of foundational model is abundant.
+Inspired by this, here we present RO-LMM, a multi-purpose, comprehensive large multimodal model (LMM) tailored for the field of radiation oncology.
+This model effectively manages a series of tasks within the clinical workflow, including clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation by leveraging the capabilities of LMM.
+In particular, to perform consecutive clinical tasks without error accumulation, we present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the consistency of handling clean inputs.
+We further extend this concept to LMM-driven segmentation framework, leading to a novel Consistency Embedding Segmentation (CESEG) techniques.
+Experimental results including multi-centre validation confirm that our RO-LMM with CEFTune and CESEG results in promising performance for multiple clinical tasks with generalization capabilities.
+Index Terms:
+Large multimodal model, Breast cancer, Clinical report, Radiotherapy target volume, Segmentation
+I Introduction
+[…]
+figureRO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology.
+The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and 3D target volume segmentation.
+Recently, the emergence of a new generation of AI models inspired by the large language models (LLMs), known as foundation models, marks a significant departure from previous paradigms [1].
+These models are characterized by their massive scale and versatility, which stems from self-supervised training on a diverse array of data.
 These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [2, 3, 4, 5, 6, 7].
@@
 Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [8].
-Specifically, in contrast to these unimodal AIs, generalist medical AIs can encompass a holistic understanding of clinical workflows, which can receive and reason a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [41, 46, 39, 50, 34].
-Inspired by this paradigm shift in medical AIs, here we present RO-LMM, a prototype of a multi-purpose medical large multimodal model (LMM), tailored to support the clinical workflow in radiation oncology.
-RO-LMM is adept at performing a variety of clinical tasks: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes.
+Specifically, in contrast to these unimodal AIs, generalist medical AIs that synergistically combine foundation models can encompass a holistic understanding of clinical workflows, which can receive a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [9, 10, 11, 1, 12].
+By understanding diverse data types and their interrelationships, multimodal AIs would facilitate more accurate diagnoses, personalized treatment development, and a reduction in medical errors by providing a comprehensive view of patient data.
+In the field of radiation oncology, which the main focus of this article, the integration of multiple modalities holds great importance so that it is one of the most important clinical fields to evaluate the potential of foundation models.
+Accordingly, here we introduce RO-LMM, a prototype large multimodal model (LMM) specifically designed to support the clinical workflow in radiation oncology.
+In particular, this work significantly extends our related prior work called LLMSeg[13], which focused on multimodal segmentation.
+More specifically, RO-LMM enlarges the LLMSeg’s scope by tackling a broader range of clinical tasks in radiation oncology: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes.
 Additionally, it is capable of (2) proposing appropriate treatment plans from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed treatment plans.
-This multifaceted functionality demonstrates a significant alignment with the expertise of clinical professionals.
-To the best of our knowledge, our prototype for assisting in the radiation oncology workflow is the first work.
-As for an initial trial, we opted for a phased approach, starting with high prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging, i.e., breast cancer.
-In the course of the aforementioned sequential tasks, the inevitable accumulation of errors in generations arises on both treatment plan suggestion and target volume segmentation task.
-To address this issue, we firstly explore Noisy Embedding Fine-Tuning (NEFTune) [19] which involves injecting uniform noise into embeddings during the training for our targeted tasks.
-To further enhance the model’s applicability, we introduce a novel consistency regularization method, bolstered by our pioneering Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs.
-Expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG).
-These advancements collectively contribute to a marked improvement in the model’s generalization capabilities in both internal and external validation.
+This multifaceted functionality of RO-LLM demonstrates a significant advances toward the expertise of clinical professionals.
+In training LLM to perform these sequential tasks from treatment plan suggestion to target volume segmentation, we found that there exist potentials for error accumulation from each tasks which can potential leads to significant end-to-end performance loss.
+Accordingly, another important contribution of this work is the employment and extension of a Noisy Embedding Fine-Tuning (NEFTune) [14] which involves injecting uniform noise into embeddings during the training for each targeted task.
+More specifically, to further enhance the model’s applicability, we develop a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs.
+Furthermore, by expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in novel Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG).
+These advancements prevent from error propagation from each subsequent task and collectively contribute to a marked improvement in the end-to-end model’s generalization capabilities in both internal and external validation.
+As for the proof of the concept study, our RO-LLM framework was adopted for breast cancer study, one of the high-prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging.
 Our contributions can be summarized as:
-  • We propose a comprehensive framework, denoted as RO-LMM, wherein LMM assists the entire workflow of radiation oncology for breast cancer treatment.
-  • We explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG.
-  • Through experiments on both internal and external datasets on real clinical breast cancer data, we demonstrated that RO-LMM outperforms baseline methods.
-2 Related Works
-2.1 Instruction Fine-tuning
-Instruction fine-tuning has been emerged as a pivotal technique for augmenting model responsiveness in the field of natural language processing (NLP).
+  • We propose a comprehensive framework, dubbed as RO-LMM, wherein LMM assists the broad workflow of radiation oncology for breast cancer treatment.
+To the best of our knowledge, our prototype is the first to assist in the comprehensive workflow of radiation oncology.
+  • To prevent from potential error accumulation throughout consecutive clinical tasks such as clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation, we explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG that significantly enhance the robustness of our method.
+  • Through experiments on various validation settings with real clinical data of breast cancer patients, we demonstrated that RO-LMM outperforms baseline methods.
+II Related Works
+II-A Instruction Fine-tuning in LLM
+In the field of LLM, the instruction fine-tuning has been emerged as a pivotal technique for augmenting model responsiveness.
 The simplicity and effectiveness of this approach have been introduced in numerous works, with a notable scale-up facilitated by advancements of large language models (LLMs) such like GPT-3 [15], ChatGPT [16], and GPT-4 [17].
 Notably, a pioneering contribution, Self-Instruct [18], involves fine-tuning of foundation models using instruction-output pairs generated from InstructGPT [19].
-This methodology, along with similar approaches, has led to diverse language model variants, including Alpaca [42], Vicuna [4], Dolly [8], and a series of LLaMA [43, 44], by exhibiting promising performance across a wide range of tasks.
+This methodology, along with similar approaches, has led to diverse language model variants, including Alpaca [20], Vicuna [21], Dolly [22], and LLaMA2 [23], by exhibiting promising performance across a wide range of tasks.
 In the medical domain, Chat-Doctor [24], Med-Alpaca [25], PMC-LLaMA [26], and Asclepius [27] have been fine-tuned for clinical question-answering (QA) task.
 Although these models demonstrate robust performance for diverse QA benchmarks, there is a notable gap in their applicability for the practical clinical workflow in a specific domain, such as radiation oncology, which serves as our targeted focus.
-2.2 Stabilizing LLM Fine-tuning with Noise
+II-B Stabilizing LLM Fine-tuning with Noise
 To enhance the robustness of language models against noisy input, various strategies have been explored.
@@
 This extension aims to further improve the robustness and generalization capabilities of language models when faced with noisy input.
-Figure 2: Training schematics of the multi-purpose RO-LMM for three different tasks.
+Figure 1: Schematics of RO-LMM training for three different tasks.
 (a) RO-LMM-S for clinical note summarization.
@@
 (c) RO-LMM-SEG++ for plan-guided target volume segmentation.
-2.3 Language-driven Image Segmentation
-Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [26], open-vocabulary segmentation [11, 51, 29], referring segmentation [48], and reasoning segmentation [25].
+II-C Language-driven Image Segmentation
+Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [32], open-vocabulary segmentation [33, 34], referring segmentation [35], and reasoning segmentation [7].
 Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable.
@@
 Target volume segmentation in the field of radiation oncology is more challenging, due to its intrinsic need for considerations of the clinical aspects beyond the image, such as overall cancer stage, treatment aim, pathological findings, and so on [38, 39, 40].
-Recent research on clinical context-aware breast cancer radiotherapy delineation has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce [35].
-Unlike previous approaches, we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
-3 Methods
-In this section, we provide a detailed methodology for our proposed approach, which is designed for sequential text generation tasks, including summarization and suggestions, as well as text-driven image segmentation, as illustrated in Figure 2.
-In Sec 3.1, we describe text-related fine-tuning approaches that yield report summarization and plan suggestion, respectively.
-In Sec 3.2, we introduce a novel segmentation approach that involves injecting noise and regularizing consistency to the text embedding for improving robustness of plan-guided 3D target volume segmentation.
-3.1 Consistency Embedding Fine-tuning for Clinical LMM
+Recent research on clinical context-aware breast cancer radiotherapy delineation [13] has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce.
+By extending [13], we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
+III Methods
+In this section, we provide a detailed description of our proposed approach designed for sequential text generation tasks, including summarization and suggestions, as well as text-driven image segmentation, whose robustness is improved by consistency embedding finetuning.
+The overall framework is illustrated in Figure 1.
+III-A Consistency Embedding Fine-tuning for Clinical LMM
 To realize the multi-purpose LMM with expertise in clinical report summarization and radiation treatment plan suggestion, we conduct instruction fine-tuning for LLaMA2 [23].
-Considering the intended objective of each task, we adopt separate strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert).
-We train a summary expert using collected notes.
+Considering the nuanced differences in the intended objective of each task, we adopt separate training strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert).
+Specifically, we train a summary expert using collected raw clinical report and summary notes.
 During inference, the summary expert receives raw clinical reports, just like in the training scenario.
 However, for the plan expert, there is a discrepancy between the training and inference scenarios.
-In other words, we opt to use the training set made up of collected notes instead of generated notes, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure 1.
-In contrast to training, our model takes the generated notes as input for inference.
-To enhance robustness for the noisy input, Noisy Embedding Fine-Tuning (NEFTune) [19] which inject uniform noise into embedding is also an effective naive solution to handle noisy inputs in this task.
-However, a crucial consideration arises from the nature of the generated notes, which may lie closer to clean inputs (collected notes) or deviate towards noisy inputs.
+In other words, we use the training set made up of collected summary notes instead of generated notes from the summary expert, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure I.
+However, at the inference phase, our model takes the generated notes from the trained summary expert.
+To deal with the input domain differences for the training and inference time, Noisy Embedding Fine-Tuning (NEFTune) [14] which inject uniform noise into embedding could be an effective naive solution to handle noisy inputs in this task.
+However, a crucial consideration arises from the nature of the generated notes, since some of them may lie closer to clean inputs (collected notes) and the others deviate towards noisy inputs.
 To address this, it is essential to train the model to handle both clean and noisy inputs.
-To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++†$\dagger$†$\dagger$†$\dagger$+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively.
+To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++111‘+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively.
 More details are as follows.
-Revisiting Noise Embedding Fine-Tuning.
-NEFTune has been recently introduced to enhance the performance of instruction fine-tuning.
-This approach involves injecting a random noise vector into embeddings during the training process as follows:
+Consistency Embedding Fine-Tuning. To enhance the performance of instruction fine-tuning, NEFTune  [14] involves injecting a random noise vector into embeddings during the training process as follows:
 $\mathcal{L}_{\text{NEFTune}}(\theta)=\mathbb{E}_{{({\mathbf{x}^{\text{emb}}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})\sim D}}\mathcal{L}_{\text{ce}}(f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}),{\mathbf{y}_{\text{txt}}}),$
-${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}={\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\boldsymbol{\epsilon,\quad\epsilon}\sim\mathcal{U}(-1,1)$ (1)
+${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}={\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\bm{\epsilon,\quad\epsilon}\sim\mathcal{U}(-1,1)$ (1)
 where ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$ is the embedding of data, $B$ denotes batch size, $L$ is token length, $C$ is embedding dimension, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is perturbed embedding, $f_{\theta}(\cdot)$ represents the model parameterized by $\theta$, ${\mathbf{y}_{\text{txt}}}$ is the label of text sample.
 They demonstrated that the effectiveness of incorporating a random noise vector adjusted based on token length, which yields robust results in most fine-tuning scenarios.
-Consistency Embedding Fine-Tuning.
+Figure 2: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module.
+These module outputs are aligned through (c) multimodal alignment module.
 To bolster the model’s capability in handling both noisy and clean inputs for treatment plan suggestion task, we incorporate a regularization loss to encourage consistency as follows:
@@
 where $\theta^{-}$ represents the model with stopped gradient, and $d(\cdot,\cdot)$ is used to quantify the discrepancy between $f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ and $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$.
-The function $\mathcal{T}$ serves as either the identity function or performs detokenization to generate sentences.
-Our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs.
 The metric function $d$ can be any capable of measuring distance in a specific space.
-However, solely minimizing distance in the same embedding space might weaken the robustness effect and pose limitations on enforcing semantic similarity.
-In our approach, to capture textual similarity between sentences, we calculate the distance in the feature space of SentenceBERT [40].
-i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [40], $\mathcal{T}$ is set to detokenization operator.
-We provide algorithmic descriptions of CEFT in Supplementary Material.
-3.2 LMM-assisted 3D Target Volume Segmentation
-For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LMM for context-aware 3D segmentation framework introduced in [35].
-However, as the LMM-assisted 3D segmentation model gets the generated treatment plan as text conditions which are inevitably noisy from the previous summarization step, we introduce noise augmentation and consistency regularization modules for the target volume segmentation network, namely, Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG).
+The function $\mathcal{T}$ performs detokenization to generate sentences.
+By introducing the consistency in the detokenized domain through $\mathcal{T}$, our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs.
+Since solely minimizing distance in the same embedding space may weaken the robustness and pose limitations on enforcing semantic similarity, to capture textual similarity between sentences, another contribution of our work is to calculate the distance in the feature space of SentenceBERT [41].
+i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [41].
+III-B LMM-assisted 3D Target Volume Segmentation
+For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LLM for context-aware 3D segmentation introduced in LLMSeg[13].
+LLMSeg demonstrated significant performance improvement in multimodal radiotherapy target segmentation compared to unimodal models.
+Additionally, it showed the ability to align target contours with provided clinical text information.
+However, in the extended clinical workflow proposed here, the LMM-assisted segmentation model receives noisy text descriptions of the generated treatment plan from the previous summarization step.
+To address this noise, we introduce two novel modules: Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG).
+These modules enhance the target volume segmentation network by handling noise and promoting consistency.
+The details are as follows.
 Noise Embedding Segmentation (NESEG).
-Firstly, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, and denote it as NESEG.
-The loss function for NESEG can be formulated as:
+Firstly, we introduce the noise augmentation technique, dubbed as NESEG, for the target volume segmentation task.
+We refer to the corresponding RO-LLM as RO-LMM-SEG+.
+Specifically, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, leading to the follow loss function:
 $\mathcal{L}_{\text{NESEG}}(\Theta)$ $=\mathbb{E}_{({\mathbf{x}_{\text{img}}},{\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}},{\mathbf{y}_{\text{img}}})\sim D}\mathcal{L}_{\text{ce}}(g_{\psi}({\mathbf{x}_{\text{img}}};\mathbf{{\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}}),{\mathbf{y}_{\text{img}}})$
@@
 where $\Theta$ = [$\psi,\phi$], ${\mathbf{x}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D CT scan, $B$ denotes batch size, $H$, $W$, and $S$ correspond to height, width, and slice of the CT scan, ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ is the perturbed text embedding, ${\mathbf{y}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D ground truth segmentation mask, $g_{\psi}(\cdot)$ represents the multimodal segmentation model parameterized by $\psi$, $f_{\theta^{*}}$ denotes a frozen LLM, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is the noise injected plan embeddings, and $\mathbf{z}_{\phi}$ is the learnable text prompt parametreized by $\phi$.
-We denote this model as RO-LMM-SEG+.
 Consistency Embedding Segmentation (CESEG).
-We further adapt the consistency regularization module for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan, and denote it as CESEG.
-Apart from the original concept of CEFTune, we modified the consistency regularization module for the multimodal segmentation task, which combines CEFTune and the text prompt tuning.
-Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is regularized with that of clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$.
-The loss function for CESEG is formulated as:
+We further adapt the consistency regularization module, dubbed as CESEG, for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan.
+The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 2.
+Apart from the original concept of CEFTune, CESEG modifies the consistency regularization module for the multimodal segmentation task, by combining CEFTune and the text prompt tuning.
+Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is also regularized with respect to the clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$.
+Specifically, the loss function for CESEG is formulated as:
 $\mathcal{L}_{\text{CESEG}}(\Theta)$ $=$ $\mathcal{L}_{\text{NESEG}}(\Theta)+\lambda\mathcal{R}(\phi,\phi^{-})$ (4)
@@
 where, $\phi^{-}$ represents the learnable text prompts with stopped gradient, $d(\cdot,\cdot)$ denotes the cosine similarity, and ${\mathbf{y}^{\text{emb}}_{\text{txt}}}=f_{\theta^{*}}({\mathbf{x}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi^{-}})\in\mathbb{R}^{B\times 1\times C}$ is the text embedding for the clean plan ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$.
-The algorithm for CESEG is provided in Supplementary Material.
-Figure 3: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module.
-These module outputs are aligned through (c) multimodal alignment module.
-The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 3.
 We employ the 3D Residual U-Net [42] as for the backbone structure within the image module, as shown in (a).
@@
 Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module.
-Further details of each module are deferred to Supplementary Material.
-4 Experiments
-4.1 Dataset
+IV Experiments
+IV-A Dataset
 To train our model, we collected internal data cohort composed of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center.
@@
 To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital.
-The detail of dataset is described in Table 1 and Supplementary Material.
+The detail of dataset is described in Table I and Supplementary Material.
 We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals.
 This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
-Table 1: Details of data for training our multi-purpose RO-LMM. “US" denotes Ultrasound. “Path" indicates Pathology.
-Task Input Output Internal Dataset External Dataset
+IV-B Implementation details
+For instruction fine-tuning for RO-LMM-S and RO-LMM-P, we use the LLaMA-2-7B-Chat [23] model as a baseline.
+The maximum context length is set at 4096, and the batch size is set to 2.
+To train RO-LMM-SEG, we employ the 3D Residual U-Net [42] using the open-source library MONAI [^2] https://monai.io/.
+For incorporating the textual information within the multimodel segmentation model, we use the LLaMA-2-7B-Chat model initialized with the pre-trained checkpoint of RO-LMM-P++.
+We set the batch size as 1.
+We use the AdamW [46] optimizer for all the tasks, with a learning rate of 5e-5 until reaching 3 epochs for both RO-LMM-S and RO-LMM-P, and 1e-4 until reaching 100 epochs for RO-LMM-SEG, leveraging 4 NVIDIA A6000 GPUs for each task.
+The hyper-parameter $\lambda$ for CEFTune is set to 1 in all of tasks, and for consistency regularization we adopt the variants of SBERT [41] which is trained on PubMed [^3] https://www.ncbi.nlm.nih.gov/pubmed/ dataset, called PubMedBERT.
+[^4] https://huggingface.co/NeuML/pubmedbert-base-embeddings.
+TABLE I: Training data details. CRS: Clinical Report Summarization. TPS: Treatment Plan Suggestion. PTS: Plan-guided Target Segmentation. US: Ultrasound. Path: Pathology.
+Task Input Output Internal External
 Train Validation Validation
-Clinical Report Summarization MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
-Treatment Plan Suggestion Clinical Note Treatment Plan
-Plan-guided Target Segmentation Treatment Plan &\&& CT Scan Segmentation 593 79 81
-4.2 Implementation details
-[…]
-[…]
-4.3 Evaluation Metric
-The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [30].
-However, acknowledging potential limitations in domain specificity for the treatment plan suggestion task, we conduct expertise-based evaluations involving two clinical experts and GPT-4.0.
-A board-certified expert curates a total of 5 scoring rubrics for evaluation, with detailed information available in Table S9.
-Using the same scoring rubrics, two clinical experts assess the performance on the plan suggestion task against ground truth labels, assigning a score of 0 or 1 for each rubric separately.
-Inspired by the recent evaluation methodologies such as GPT-Score [13] and G-Eval [32], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning.
-To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [9] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
-To further validate effectiveness of our experiment, we provide the confidence intervals (CI) on each metric by utilizing the non-parametric bootstrap method.
-We calculate CI by conducting 1,000 times of random sampling procedure with replacement from the experimental results, each time equal to the total size of the sample dataset.
-Subsequently, we estimate the 95th percentile CI from the relative frequency distribution of each experimental trial.
-4.4 Baseline Models
-For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [14], ChatDoctor [52], Asclepius [24], and PMC-LLAMA [49].
-Additionally, we include a comparison with ChatGPT [36] utilizing few-shot in-context learning for both summary and treatment plan suggestion task.
-For the segmentation task, we utilize 3D Residual U-Net [7] and ConTEXTualNET [18] as our baseline methods.
-We further utilize an additional baseline as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
-Table 2: The quantitative results for clinical note summarization using different methods. Vanilla denotes the instruction fine tuning. CI denotes confidence interval.
+CRS MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
+TPS Clinical Note Treatment Plan
+PTS Treatment Plan &\&& CT Scan Segmentation 593 79 81
+TABLE II: Quantitative comparison for clinical note summarization. Vanilla: the instruction fine tuning. CI: confidence interval.
 Model Method Internal Validation (N=60) External Validation (N=81)
@@
 PMC-LLaMA [26] - 0.118 (0.101-0.137) 0.034 (0.029-0.040) 0.118 (0.101-0.137) 0.165 (0.136-0.195) 0.050 (0.040-0.060) 0.165 (0.136-0.195)
-ChatGPT - 0.453 (0.412-0.488) 0.215 (0.187-0.246) 0.453 (0.412-0.488) 0.390 (0.357-0.423) 0.193 (0.171-0.215) 0.390 (0.357-0.423)
+GPT4.0 - 0.366 (0.343-0.391) 0.230 (0.207-0.252) 0.366 (0.343-0.391) 0.316 (0.294-0.339) 0.249 (0.226-0.272) 0.316 (0.294-0.339)
 RO-LMM-P Vanila 0.651 (0.617-0.689) 0.486 (0.437-0.539) 0.651 (0.617-0.689) 0.591 (0.574-0.609) 0.426 (0.403-0.452) 0.591 (0.574-0.609)
@@
 RO-LMM-P++ CEFT 0.668 (0.632-0.705) 0.512 (0.460-0.564) 0.668 (0.632-0.705) 0.615 (0.592-0.638) 0.459 (0.429-0.489) 0.615 (0.592-0.638)
-4.5 Results
-Clinical Report Summarization.
-We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table 2.
-Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals.
-Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning.
-We also perform expert analysis on the summarized reports, and the results and experimental details are deferred to Supplementary Material.
-The qualitative assessment in Table 4 further highlights the effectiveness of our proposed model, providing well-organized content and consistent formatting compared to the ground truth label.
-While there were minor errors observed in terms of the specific distance measurement from the nipple (red), which appears to stem from the process of integrating ultrasound and MRI images, our model accurately summarizes information in a format consistent with the provided labels.
-In contrast, Chat-GPT exhibits issues such as omitting essential information and generating hallucinated content (red).
-Treatment Plan Suggestion.
-As indicated in Table 2, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes.
-Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines.
-Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach.
-However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
-Furthermore, for the treatment plan suggestion task, we comprehensively evaluate the generated plans using our proposed expertise-based rubrics as shown in Table S9.
-Table 3 presents the clinical expert analysis results, where R##\## represents each rubric and C##\## represents each clinical expert.
-For both internal and external validation, our RO-LMM variants significantly outperform the results of GPT4.0 with few-shot in-context learning due to the specific knowledge of the domain.
-We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals.
-We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$).
-The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
-For a more detailed understanding, examples of the treatment plan suggestion task are provided in Table 4.
-The proposed RO-LMM-P++ (Ours) consistently yields correct answers compared to the labeled ground truth.
-These results demonstrate the effectiveness of our proposed CEFTune for this specific task.
-While Chat-GPT has shown success in providing some level of information regarding where to treat, it falls short when it comes to suggesting the specific treatment regions and dose scheme (red).
-Table 3: Clinical expert analysis for treatment plan suggestion. R# represents each rubric and C# represents each clinical expert.
+TABLE III: Clinical expert analysis for report summarization. R#: each rubric, C#:each clinical expert.
+Internal Validation (N=10)
+ChatGPT LLaMa2-7B RO-LMM-S
+Experts R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total $r$
+C1 3.0 / 0.0 / 4.0 / 7.0 / 13.3 2.0 / 0.0 / 0.0 / 7.0 / 9.0 15.0 / 10.0 / 10.0 / 10.0 / 45.0 0.884
+C2 3.0 / 0.0 / 5.0 / 7.0 / 14.4 1.0 / 0.0 / 0.0 / 9.0 / 10.0 15.0 / 10.0 / 9.0 / 10.0 / 44.0
+External Validation (N=10) $r$
+C1 5.0 / 1.0 / 9.0 / 7.0 / 22.0 7.0 / 0.0 / 0.0 / 7.0 / 14.0 10.0 / 9.0 / 8.0 / 8.0 / 35.0 0.977
+C2 6.0 / 3.0 / 9.0 / 7.7 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
+IV-C Evaluation Metric and Baseline Models
+The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [47].
+However, acknowledging potential limitations in domain specificity for both tasks, we conduct evaluations using two clinical experts and GPT-4.0.
+A board-certified expert creates five scoring rubrics for plan suggestions, detailed in Table A1.
+Using these rubrics, the experts score each aspect of the plans, assigning 0 or 1 for each rubric separately.
+Similarly, for clinical report summarization, rubrics are created and the same evaluation process is followed.
+Inspired by the recent evaluation methodologies such as GPT-Score [48] and G-Eval [49], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning on the treatment plan task.
+To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [50] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
+To further validate effectiveness of our experiment, we provide the confidence intervals (CI) on each metric by utilizing the non-parametric bootstrap method.
+We calculate CI by conducting 1,000 times of random sampling procedure with replacement from the experimental results, each time equal to the total size of the sample dataset.
+Subsequently, we estimate the 95th percentile CI from the relative frequency distribution of each experimental trial.
+For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [25], ChatDoctor [24], Asclepius [27], and PMC-LLAMA [26].
+Additionally, we include a comparison with ChatGPT [16] for the report summarization task and GPT-4.0 for the treatment plan suggestion task using few-shot in-context learning.
+For the segmentation task, we utilize 3D Residual U-Net [42] and ConTEXTualNET [37] as our baseline methods.
+We further utilize an additional baseline as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
+TABLE IV: Clinical expert analysis for treatment plan suggestion. R#: each rubric, C#: each clinical expert.
 Internal Validation (N=25)
-GPT4.0 RO-LMM-P RO-LMM-P+ (NEFT) RO-LMM-P++ (CEFT)
+GPT-4.0 RO-LMM-P RO-LMM-P+ (NEFT) RO-LMM-P++ (CEFT)
 Experts R1 / R2 / R3 / R4 / R5 / Total R1 / R2 / R3 / R4 / R5 / Total R1 / R2 / R3 / R4 / R5 / Total R1 / R2 / R3 / R4 / R5 / Total $r$
@@
 C2 9.6 / 9.6 / 5.6 / 4.4 / 3.8 / 32.7 5.6 / 10.0 / 8.4 / 8.8 / 9.2/ 42.0 5.6 / 9.6 / 8.0 / 8.4/ 10.0 / 41.6 4.4 / 10.0 / 8.8 / 9.2 / 9.6 / 42.0 0.4
-GPT 10.0 / 7.6 / 4.8 / 5.6 / 6.8 / 29.2 8.2 / 8.0 / 7.2 / 8.8 / 9.6 / 37.2 9.1 / 7.6 / 6.4 / 7.6 / 9.2 / 34.8 9.1 / 8.4 / 8.0 / 8.8 / 10.0 / 39.2
+GPT-4.0 10.0 / 7.6 / 4.8 / 5.6 / 6.8 / 29.2 8.2 / 8.0 / 7.2 / 8.8 / 9.6 / 37.2 9.1 / 7.6 / 6.4 / 7.6 / 9.2 / 34.8 9.1 / 8.4 / 8.0 / 8.8 / 10.0 / 39.2
 External Validation (N=25) $r$
@@
 C2 8.8 / 8.5 / 6.2 / 3.4 / 7.3 / 34.2 8.5 / 9.6 / 7.7 / 7.3 / 9.2 / 42.3 8.1 / 9.6 / 8.1 / 8.8 / 9.2 / 43.8 8.5 / 9.6 / 8.8 / 8.8 / 9.6 / 45.8 0.4
-GPT 8.8 / 6.9 / 6.2 / 5.8 / 4.2 / 31.9 8.1 / 7.3 / 6.2 / 1.5 / 8.8 / 31.9 8.5 / 8.1 / 5.0 / 1.2 / 9.2 / 31.9 8.5 / 8.1 / 6.5 / 3.8 / 9.2 / 36.2
-Table 4: Example case for the clinical report summarization task and plan suggestion task compard to baseline methods.
-The red scripts indicate incorrect or redundant information, and inconsistent formatting compared to the label, evaluated by a board certified clinical expert.
-Clinical Report Summarization Example
-User
-Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
-##\## Input: [MRI reports ]…[Pathology reports] …
-Ground Truth
-[…]
-Ours
-[…]
-Chat-GPT
-[…]
-Treatment Plan Suggestion Example
-User
-Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
-##\## Input: [Clinical notes]…
-Ground Truth
-RT to Lt. whole breast + IMN + SCL + AXL 40.05 Gy / 15 fx $\rightarrow$ 10 Gy boost (or 48 Gy / 15 fx with IMRT).
-Ours
-RT to Lt. whole breast, IMN, SCL, AXL 40.05 Gy / 15 fx + 10 Gy boost (IMRT: 48 Gy / 15 fx).
-Med-Alpaca
-..Breast, LLO 4:00 3cm FN, IDC, icT2N1M0, stage IIB $\rightarrow$ ypT1cN1M0, stage IIA…
-ChatDoctor
-…The T-stage, N-stage, ER/PR/HER2/Ki67, and Lum A status of the tumor, as well as the patient’s chemotherapy and surgical history, can all be used to determine appropriate radiation …
-Chat-GPT
-RT to Left breast, chest wall, L axilla Conventional fractionation 40.05 Gy / 15 fx.
-Table 5: Comparison of 3D target volume segmentation performance.
+GPT-4.0 8.8 / 6.9 / 6.2 / 5.8 / 4.2 / 31.9 8.1 / 7.3 / 6.2 / 1.5 / 8.8 / 31.9 8.5 / 8.1 / 5.0 / 1.2 / 9.2 / 31.9 8.5 / 8.1 / 6.5 / 3.8 / 9.2 / 36.2
+TABLE V: Comparison of 3D target volume segmentation performance.
 Model LMM Tuning Internal Validation (N=79) External Validation (N=81)
@@
 RO-LMM-SEG++ CESEG 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
-${}^{\dagger}$modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
-Table 6: Comparison of 3D target segmentation performance between unimodal and multimodal models for both overall and specific patient types.
+modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
+TABLE VI: Comparison of 3D target segmentation performance for overall and specific patient types.
 Types Method Internal Validation External Validation
@@
 Gain $\uparrow$ +0.040 +0.044 -5.571 +0.215 +0.199 -4.728
+IV-D Results
+Clinical Report Summarization.
+We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table II.
+Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals.
+Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning.
+Moreover, we evaluate the generated summaries using expertise-based rubrics by two clinical experts and compare them to baselines, including ChatGPT and LLaMa-2.
+As shown in Table III, our ROB-LMM-S model significantly outperforms all baselines in both internal and external validations, thanks to its domain-specific knowledge.
+Additionally, Pearson correlation ($r$) analysis reveals strong positive inter-clinician correlations ( $>0.85$ and $>0.95$ for internal and external validation, respectively), confirming the reliability of our rubrics and the clinical relevance of ROB-LMM-S.
+Therefore, our ROB-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
+Treatment Plan Suggestion.
+As indicated in Table II, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes.
+Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines.
+Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach.
+However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
+Furthermore, we comprehensively evaluate the generated plans using our expertise-based rubrics and clinical expert analysis, as detailed in Table IV.
+For both internal and external validation, our RO-LMM variants significantly outperform GPT-4.0 with few-shot in-context learning, as assessed by both GPT and clinical experts, due to their domain-specific knowledge.
+We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals.
+We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$).
+The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
 3D Target Volume Segmentation.
-As reported in Table 5, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting.
+As reported in Table V, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting.
 This can be also observed in qualitative results.
 As shown in Figure 3(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks.
-Moreover, as shown in Figure 4(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows.
-RO-LMM-SEG without any proposed module shows reasonable performance, yet sometimes undersegments the treatment target volume, as shown in Figure 4(b).
+Moreover, as shown in Figure 3(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows, and RO-LMM-SEG sometimes under-segments the target volume.
 In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated.
 We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type).
-The comparative results are presented in the Table 6 separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG.
+The comparative results are presented in the Table VI separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG.
 Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively.
@@
 This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
-Figure 4: Qualitative comparison on 3D target volume segmentation task between unimodal and multimodal approaches. Red arrows indicate incorrectly segmented regions.
-5 Discussion
-5.1 Analysis on Strategy of Separate Expert
-To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing a unified model shown in Table 8.
-We observe that the unified model’s performance is significantly inferior to our distinct training strategy.
-This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task.
-These results suggest that the knowledge required for each summarization and suggestion task necessitates task-aware training strategies.
-Table 7: Ablation study on adopting separate expertise for each textual task against unified strategy.
-Table 8: Ablation study on CESEG with input text variation for target segmentation performance.
+Inference on Public Dataset.
+To validate the effectiveness of our methods when AI-generated results are reused as input, we evaluated the performance by inputting each model’s report summarization results its plan suggestion task.
+We conducted experiments on a publicly available text-related dataset, as detailed in Table VII.
+Our proposed methods consistently outperformed other specialized medical LLMs and variants of GPT (such as ChatGPT and GPT-4) in both clinical report summarization and therapy plan suggestion tasks.
+Notably, our RO-LMM-P++ method demonstrated substantial performance enhancements, with improvements of up to 0.55 and 0.31 over other medical LLMs and GPT variants, respectively, demonstrating our model’s robustness to error accumulation during sequential generations.
+TABLE VII: Quantitative comparison results for our RO-LMM’s clinical report summarization and therapy plan suggestion performance on the publicly available dataset.
+Model Pubilc available dataset (N=15)
+Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$ (CI) Rouge L $\uparrow$ (CI)
+Clinical Report Summarization
+MedAlpaca [25] 0.597 (0.502-0.683) 0.431 (0.343-0.516) 0.597 (0.502-0.683)
+ChatDoctor [24] 0.644 (0.511-0.757) 0.476 (0.370-0.573) 0.644 (0.511-0.757)
+Asclepius [27] 0.696 (0.658-0.733) 0.469 (0.444-0.493) 0.696 (0.658-0.733)
+ChatGPT 0.765 (0.733-0.795) 0.526 (0.498-0.554) 0.765 (0.733-0.795)
+Ours (RO-LMM-S) 0.795 (0.746-0.833) 0.569 (0.530-0.612) 0.795 (0.746-0.833)
+Therapy Plan Suggestion
+MedAlpaca [25] 0.165 (0.111-0.233) 0.085 (0.064-0.109) 0.165 (0.111-0.233)
+ChatDoctor [24] 0.189 (0.147-0.233) 0.104 (0.085-0.122) 0.189 (0.147-0.233)
+Asclepius [27] 0.125 (0.102-0.151) 0.074 (0.061-0.089) 0.125 (0.102-0.151)
+GPT-4.0 0.390 (0.302-0.485) 0.231 (0.179-0.292) 0.390 (0.302-0.485)
+Ours (RO-LMM-P++) 0.669 (0.588-0.740) 0.450 (0.374-0.525) 0.669 (0.588-0.740)
+Figure 3: Qualitative comparison on 3D target volume segmentation task. Red arrows indicate errors.
+TABLE VIII: Ablation study on adopting separate expertise for each textual task against unified strategy.
 Training Strategy Internal (N=60) External (N=81)
@@
 Distinct strategy 0.652 / 0.491 / 0.649 0.606 / 0.450 / 0.606
+TABLE IX: Ablation study on CESEG for target segmentation performance with input text variation.
 CESEG Input Text Target Segmentation
@@
 Difference $\downarrow$ 0.007 0.010
-Table 8: Ablation study on CESEG with input text variation for target segmentation performance.
-5.2 Analysis on Consistency Regularization
+V Ablation Study
+Analysis on Strategy of Separate Expert To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing another baseline model that unifies summarization and plan suggestion, and show the results in Table IX.
+We observe that the unified model’s performance is significantly inferior to our distinct training strategy.
+This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task.
+These results suggest that the knowledge required for each task necessitates task-aware training strategies.
 […]
-5.3 Limitation
-While our study provides valuable insights on the application of radiation oncology, there are several limitations to consider.
-Firstly, the dataset used in this study is currently confined to patients with their initial diagnoses, primarily focusing on breast cancer.
-Expanding the scope of our research to include diverse patient scenarios and other cancer types is essential for enhancing the applicability of our model.
-Secondly, achieving detailed target volume segmentation for Simultaneous Integrated Boost (SIB) necessitates information beyond treatment plan-guided segmentation.
-This additional information may be obtained from alternative imaging modalities, such as MRI or PET-CT, along with ground truth masks for organ-at-risks (OARs) and gross tumor volume (GTV).
-Consequently, SIB is excluded from this study.
-Future research will focus on developing a model for SIB by incorporating a more extensive array of information.
-6 Conclusion
-In this work, we introduce RO-LMM, a multi-purpose foundation model tailored for radiation oncology.
-Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation, mirroring real-world clinical workflows.
+TABLE X: Inference computational complexity. RM: RO-LMM.
+External Validation (N=81)
+RM-S RM-P++ RM-SEG++ Analysis
+Memory (GB) 13.68 13.57 17.89 Mean 14.93
+Duration (sec/patient) 6.89 $\pm$ 1.68 1.75 $\pm$ 0.51 1.54 $\pm$ 0.72 Total 10.19
+VI Discussion
+While our study provides valuable insights into the application of radiation oncology, several limitations must be acknowledged.
+Firstly, the dataset is confined to patients with initial breast cancer diagnoses.
+Expanding to diverse patient scenarios and other cancer types is crucial for broader applicability.
+Secondly, radiotherapy for breast cancer in real clinical setting often requires techniques like sequential boost or simultaneous integrated boost (SIB) to deliver higher radiation doses around the tumor bed.
+However, accurate tumor bed delineation requires integrating information from multiple imaging modalities, which simple plan-guided segmentation alone cannot achieve.
+Therefore, we did not include tumor bed delineation for boost techniques in this study to simplify the problem.
+Future research should focus on developing a framework capable of delineating these boost volumes to aid in clinical treatment.
+Finally, our model did not utilize all clinical information from the actual EMR for decision-making due to token length limitations.
+We only used MRI, US, and pathology results, considered most critical for initial diagnosis and treatment planning.
+Overcoming token length constraints to integrate all EMR data will be essential for future studies.
+Despite limitations, our RO-LMM demonstrates the potential to efficiently handle a broad range of clinical workflows using a single NVIDIA A6000 GPU within a 24GB memory, completing tasks in approximately 10 seconds (see Table X).
+This capability facilitates local implementation even with limited GPU resources and holds the promise of reducing the workload of radiation oncologists via the automation of clinical report-based plan suggestion for delineating radiation target volumes.
+Furthermore, our method demonstrates its effectiveness by surpassing closed-source LLMs on both a multi-center validation set and a publicly available dataset.
+These results suggest that our locally tuned RO-LMM-P++ holds significant potential to replace closed-source LLMs while maintaining data privacy and compliance with local data exposure regulations.
+VII Conclusion
+In this work, we introduce RO-LMM, a multi-purpose, comprehensive foundation model tailored for radiation oncology.
+Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks encompassing overall workflow of radiation oncology: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation.
 Another key contribution of this work is the introduction of consistency technique into both text and segmentation task.
@@
 Ethical Statement All the patient information and any potential personally identifiable information from used datasets are de-identifiied, and clinical data usage is approved by the Institutional Review Board (IRB).
-Appendix 0.A Supplementary Section
-In this supplemental documents, we present:
-  • An algorithmic descriptions of our proposed Consistency Embedding Fine Tuning (CEFT) and Consistency Embedding Segmentation (CESEG) in Section 0.B.
-  • An explanation of the network architecture for the segmentation model in Section 0.C.
-  • Results of ablation studies for both text-related tasks and target volume segmentation tasks in Section 0.D and Section 0.C.
-  • Detailed information about the dataset, including pre-processing and acquisition procedures, in Section 0.F.
-  • Examples of the entire workflow in Section 0.I.
-  • Additional comparison results for all tasks in Section 0.G.
-  • Detailed descriptions of the evaluation for treatment plan suggestion, including prompts for evaluation, in Section 0.J.
-  • Detailed descriptions of the expert-analysis for clinical report summarization, including prompts for evaluation, in Section 0.H.
-Appendix 0.B Algorithm of CEFTune and CESEG
-Algorithm 1 and Algorithm 2 details the training procedure with CEFTune and CESEG, respectively.
-Input :  $\{X^{\text{txt}}_{i},Y^{\text{txt}}_{i}\}_{i=1}^{N}$ $\sim D$ tokenized dataset, large language model $f_{\theta}(\cdot)$, detokenization operator $\mathcal{T(\cdot)}$, the metric which calculate the cosine similiary $d(\cdot,\cdot)$, the hyper-parameter for the noise intensity and loss, $\alpha$, and $\lambda$, repectively.
-The cross entropy loss $\mathcal{L}_{\text{ce}}$,
-the embedding layer $\text{emb}(\cdot)$, $\theta^{-}$leftarrow stopgrad $(\theta)$;
-1 repeat
-2       $\text{Sample}\;({\mathbf{x}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})$ $\sim D$ ;
-3       ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\leftarrow\text{emb}({\mathbf{x}_{\text{txt}}})$ ;
-4       ${\mathbf{\hat{y}}_{\text{txt}}}\leftarrow f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$ ;
-5       $\epsilon\sim\mathcal{U}(-1,1)$;
-6       ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}\leftarrow{\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\epsilon$ ;
-7       ${\mathbf{\tilde{y}}_{\text{txt}}}\leftarrow f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ ;
-8       $\mathcal{L}(\theta)\leftarrow$ $\mathcal{L}_{\text{ce}}({\mathbf{\hat{y}}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})+\lambda d(\mathcal{T}({\mathbf{\tilde{y}}_{\text{txt}}}),\mathcal{T}({\mathbf{\hat{y}}_{\text{txt}}}))$
-9until Stopping criteria met max iterations;
-Algorithm 1 CEFTune: Consistency Embedding Fine tuning
-Input :  $\{X_{i}^{\text{img}},X_{i}^{\text{txt}},Y_{i}^{\text{img}}\}_{i=1}^{M}$ $\sim D$, the entire image segmentation model $g_{\Theta}(\cdot)$, where $\Theta$ = [$\psi,\phi$], which consist of segmentation model $g_{\psi}$, and learnable text prompt $z_{\phi}$.
-The frozen large language model $f_{\theta^{*}}(\cdot)$, output embedding of $f_{\theta^{*}}$ ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$, the hyper-parameter for the noise intensity and loss $\alpha$ and $\lambda$, respectively, the metric which calculate the cosine similarity $d(\cdot,\cdot)$, the cross entropy loss $\mathcal{L}_{\text{ce}}$,
-the embedding layer $\text{emb}(\cdot)$, $\phi^{-}$leftarrow stopgrad $(\phi)$,
-1 repeat
-2       $\text{Sample}\;({\mathbf{x}_{\text{img}}},{\mathbf{x}_{\text{txt}}},{\mathbf{y}_{\text{img}}})$ $\sim D$ ;
-3       ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\leftarrow\text{emb}({\mathbf{x}_{\text{txt}}})$ ;
-4       ${\mathbf{y}^{\text{emb}}_{\text{txt}}}\leftarrow f_{\theta^{*}}({\mathbf{x}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi})$ ;
-5       $\epsilon\sim\mathcal{U}(-1,1)$;
-6       ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}\leftarrow{\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\epsilon$ ;
-7       ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\leftarrow f_{\theta^{*}}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi^{-}})$ ;
-8       $\mathcal{L}(\Theta)\leftarrow$ $\mathcal{L}_{\text{ce}}(g_{\psi}({\mathbf{x}_{\text{img}}};\mathbf{{\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}}),{\mathbf{y}_{\text{img}}})+\lambda d({\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}},{\mathbf{y}^{\text{emb}}_{\text{txt}}})$
-9until Stopping criteria met max iterations;
-Algorithm 2 CESEG: Consistency Embedding Segmentation
-Appendix 0.C Architecture of RO-LMM-SEG++
-The schematic of our RO-LMM-SEG++ is illustrated in Figure 3.
-For text prompt tuning of LLM, we introduce a total of $N$-text prompts $\{{{\mathbf{z}_{\phi}}^{n}|^{N}_{n=1}}\}$, where each ${\mathbf{z}_{\phi}}^{n}\in\mathbb{R}^{M\times C}$ is composed of $M$ vectors with the dimension $C$.
-These learnable vectors are randomly initialized, and then consistently prepended to each of treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$.
-Here, the final prompted text input $t$ can be formulated as follows:
-$t=\{{\mathbf{z}_{\phi}}_{1}^{n},{\mathbf{z}_{\phi}}_{2}^{n},...,{\mathbf{z}_{\phi}}_{M}^{n},{\mathbf{x}^{\text{emb}}_{\text{txt}}}\}.$ (5)
-Then, using the prompted text input $t$, the frozen LLM resulted the text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ as for the last token output.
-Then, to align the text embedding with the intermediate image embedding ${h}_{img}^{l}\in\mathbb{R}^{B\times H_{l}W_{l}S_{l}\times Ch_{l}}$, ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ is projected to yield ${h}_{txt}^{l}\in\mathbb{R}^{B\times 1\times Ch_{l}}$ via $l$-th layer-wise linear layer, where $H_{l}$, $W_{l}$, and $S_{l}$ correspond to height, width, and slice of the image embedding, and $Ch_{l}$ is the channel dimension of $l$-th skip-connection layer within the image module.
-For multimodal alignment, the projected text embeddings ${h}_{txt}^{l}$ is self-attended and crossly-attended with the image embedding ${h}_{img}^{l}$ for each layer, and each multimodal embedding ${h^{*}}^{l}_{img}$ is inputted to the decoder part of the image module to yield final prediction $\hat{y}_{\text{img}}\in\mathbb{R}^{B\times 1\times HWS}$.
-Table S1: Component analysis of RO-LMM-P framework on treatment plan suggestion.
-Component Condition Ours Internal Validation (N=60) External Validation (N=81)
-Rouge 1 $\uparrow$
-Rouge 2 $\uparrow$
-Rouge L $\uparrow$
-Rouge 1 $\uparrow$
-Rouge 2 $\uparrow$
-Rouge L $\uparrow$
-Teacher Module
-Noisy teacher
-0.613
-0.453
-0.615
-0.575
-0.428
-0.571
-Clean teacher
-$\checkmark$
-0.652 0.491 0.649 0.606 0.450 0.604
-SBERT Moudle
-SBERT
-0.648
-0.484
-0.645
-0.598
-0.442
-0.594
-PubMedBERT
-$\checkmark$
-0.652 0.491 0. 649 0.606 0.450 0.604
-Noise Intensity ($\alpha$) 5
-$\checkmark$
-0.652 0.491 0.649 0.606 0.450 0.604
-10
-0.648
-0.488
-0.647
-0.597
-0.435
-0.598
-15
-0.642
-0.481
-0.642
-0.576
-0.415
-0.572
-Noise type Uniform 5
-$\checkmark$
-0.652 0.491 0.649 0.606 0.450 0.604
-Gaussian 5
-0.625
-0.449
-0.629
-0.588
-0.425
-0.586
-Backbone
-LLaMA-13B
-0.645
-0.478
-0.640
-0.597
-0.436
-0.594
-+NEFT
-0.658
-0.500
-0.653
-0.601
-0.438
-0.597
-+CEFT
-$\checkmark$
-0.662 0.502 0.661 0.613 0.441 0.603
-Temperature
-0.6
-0.647
-0.480
-0.643
-0.593
-0.435
-0.591
-0.8
-$\checkmark$
-0.652 0.491 0.649 0.606 0.450 0.604
-1.0
-0.662
-0.487
-0.647
-0.601
-0.442
-0.599
-Appendix 0.D Ablation Study on Therapy Plan Suggestion
-To investigate the effectiveness of our proposed methods, we conduct the ablation studies for treatment plan suggestion task, as shown in Table S1.
-Consistency Regularization Loss We manipulate the objective function in Equation 3.1 to analyze the effect of consistency regularization.
-By default, we use $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$ as the target (referred to as “clean teacher").
-For comparison, we set $f_{\theta^{-}}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ as the teacher and $f_{\theta}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$ as the student (referred to as "noisy teacher").
-We observe that the clean teacher significantly outperforms the noisy teacher.
-By restricting the divergence of the distribution of students, the clean teacher yields enhanced performance compared to the noisy teacher.
-Effect of SBERT Module We study the effect of the SBERT module in enforcing consistency between clean and noisy input.
-Our model, equipped with the medical domain-specific SBERT variant, PubMedBERT, outperforms models trained on general language.
-This demonstrates the effectiveness of incorporating domain-specific knowledge.
-Analysis of Noise type and Intensity ($\alpha$) Since our proposed method involves injecting noise into embeddings, we conduct a study to analyze the effect of noise type and intensity under various conditions.
-First, regarding the noise type, we found that our method performed best when equipped with uniform noise compared to Gaussian noise.
-Consequently, we set the default configuration for noise type to uniform noise.
-Additionally, we analyze the effect of noise intensity by varying the parameter $\alpha$ from 5 to 15.
-Smaller values of $\alpha$ resulted in better performance on the internal dataset, while larger values improved performance on the external dataset.
-However, excessively large values of $\alpha$ led to a significant drop in performance on the external dataset.
-Based on empirical results, we select $\alpha$ set to 5 as the baseline.
-Analysis of Model Backbone We conduct experiments using the larger 13B LLaMA-2 backbone version and observ a similar performance trend compared to the 7B model.
-Despite the larger backbone size, our proposed CEFT method outperform the other baselines, demonstrating scalability.
-Taking into account both cost-effectiveness and performance, we adopt the 7B model as the default configuration.
-Analysis of Temperature To investigate the impact of stochasticity, we conducted experiments varying the temperature during treatment plan suggestion generation.
-Our findings indicate that the results produced by our proposed method are insensitive and robust to changes in temperature.
-Based on empirical analysis, we set the default temperature to 0.8.
-Appendix 0.E Ablation Study on Target Segmentation
-To optimize treatment target segmentation performance, we conduct the ablation studies, as shown in Table S2.
-Analysis of LMM Tuning Methods We compare the different LMM tuning methods to find optimal performance, where NESEG nor CESEG module is not activated.
-We analyze various methods including no tuning, low-rank adaptation (LoRA) fine-tuning [16] and prompt tuning methods with a single or multiple text prompts.
-From the experimental results, the text prompt tuning shows the most reliable performance across the external validation setting.
-Additionally, the use of single text prompt shows a remarkable improvement in performance compared to utilizing multiple text prompts.
-Analysis of Noise Intensity ($\alpha$) We analyze the effect of noise intensity by varying $\alpha$ from 5 to 20, where CESEG module is not activated.
-We empirically find that $\alpha$ as 10 yields the best performance.
-We further train the hyper-parameter $\alpha$ as a learnable parameter, however, yielding inferior performance compared to employing the fixed noise.
-Table S2: Component analysis of RO-LMM-SEG framework on target volume segmentation performance.
-Component Condition Ours Internal Validation (N=79) External Validation (N=81)
-Dice $\uparrow$
-IoU $\uparrow$
-HD-95 $\downarrow$
-Dice $\uparrow$
-IoU $\uparrow$
-HD-95 $\downarrow$
-LMM Tuning Method
-None
-$0.822$
-$0.712$
-$12.368$
-$0.752$
-$0.622$
-$24.849$
-LoRA [16]
-$0.810$
-$0.699$
-$\bf{4.989}$
-$0.759$
-$0.637$
-$\bf{16.692}$
-Text Prompts (p=1)
-$\checkmark$
-$\bf{0.825}$
-$\bf{0.714}$
-$5.392$
-$\bf{0.775}$
-$\bf{0.652}$
-$17.623$
-Text Prompts (p=2)
-$0.822$
-$0.711$
-$7.030$
-$0.757$
-$0.627$
-$24.64$
-Noise Intensity ($\alpha$)
-5
-$0.821$
-$0.707$
-$5.917$
-$0.738$
-$0.601$
-$\bf{22.320}$
-10
-$\checkmark$
-$\bf{0.830}$
-$\bf{0.720}$
-$\bf{4.383}$
-$\bf{{0.771}}$
-$\bf{{0.653}}$
-$28.835$
-20
-$0.826$
-$0.716$
-$5.575$
-$0.732$
-$0.593$
-$26.090$
-Learnable
-$0.828$
-$0.715$
-$5.024$
-$0.757$
-$0.633$
-$23.694$
-Appendix 0.F Details of Training Dataset
-For data preparation, we collected internal training & validation data from patients treated at Yonsei Cancer Center between January 2009 and December 2022.
-For external validation, we further collected data from patients from Yongin Severance Hospital between January 2018 and December 2022.
-We confirmed that the external test set was not overlapped with the training dataset.
-For pre-processing image data for training RO-LMM-SEG++, all the 3D chest CT scans and target volumes were re-sampled with identical voxel spacing of 1.0 $\times$ 1.0 $\times$ 3.0 $\text{mm}^{3}$.
-The intensity values of the CT scans were truncated between -1,000 and 1,000 of Hounsfield unit (HU), and linearly normalized between 0 and 1.0.
-When training, a 3D patch with a spatial dimension of 384 $\times$ 384 $\times$ 128 pixels was randomly cropped.
-When inference, the entire 3D CT scans was tested using sliding windows with the identical spatial dimension of the 3D patch to that used for training.
-Appendix 0.G Additional Qualitative Results
-In this section, we provide the additional qualitative results on all tasks as shown in Table S4, Table S5 and Figure S1.
-In the context of report summarization task involving cases with multiple (two or more) tumor masses, as shown in in Table S4, our model adeptly summarize the existence of two tumors in the first line of the summary, despite minor inaccuracies in pinpointing the exact locations of the tumors.
-In contrast, ChatGPT incorrectly summarize these cases as having only one tumor mass in the first line.
-Additionally, ChatGPT exhibit instances of hallucination, such as inappropriately mentioning tumor sizes in the first line, erroneously including post-therapy (yp) stages for patients who did not receive chemotherapy, and excessively condensing key information such as the presence of the lymph node metastasis, making it challenging to discern.
-Overall, unlike our model which aligns closely with the ground truth in its summaries, ChatGPT demonstrates significant issues, including the generation of erroneous hallucination and the failure to comprehensively encapsulate necessary details.
-In the evaluation of treatment plan suggestion capabilities in Table S5, only our model and ChatGPT are able to offer somewhat meaningful suggestions related to radiotherapy treatment, while other models fail.
-ChatGPT, drawing upon its built-in knowledge base, seems to propose plans that are partially relevant.
-However, it exhibitd a fundamental misunderstanding of dose schemes, frequently misinterpreting hypofractionation as conventional fractionation.
-A significant limitation in ChatGPT’s approach is its inability to incorporate all regional lymph node areas into the target volume, often disproportionately focusing on specific areas, such as the axillary lymph nodes.
-Additionally, ChatGPT struggles to distinguish between the treatment of the breast and the chest wall, leading to flawed treatment plan suggestions.
-These limitations seen to arise from ChatGPT’s lack of a holistic understanding of radiation oncology, unlike our domain-specific model.
-Figure S1 provides qualitative comparisons of the target volume segmentation performance for breast cancer patients, where the unimodal model fails to segment target volume with incorrect laterality as indicated as red arrows.
-On the other hand, our proposed multimodal RO-LMM-SEG++ with CESEG shows the most promising qualitative segmentation performance, aligning consistently with the ground truth masks.
-Figure S1: 3D target volume segmentation cases compared to baseline methods. Red arrows indicate incorrectly segmented regions.
-Appendix 0.H Expert Analysis on Clinical Report Summarization
-Similar to treatment plan suggestions, we evaluate the generated summaries using our proposed expertise-based rubrics, as illustrated in Tab.
-S9, and compare them with other baselines, including ChatGPT and LLaMa-2.
-Tab.
-S3 presents the results of clinical expert analysis, where R##\## represents each rubric and C##\## represents each clinical expert.
-For both internal and external validation, our ROB-LMM-S significantly outperforms all other baselines due to its specific knowledge in this domain.
-We further evaluate inter-clinician correlation using Pearson correlation ($r$), with results showing strong positive correlations ($>0.85$ and $>0.95$, for each internal and external validation), affirming the reliability of our proposed evaluation rubrics and the clinical effectiveness of our ROB-LMM-S.
-Therefore, our ROB-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
-Table S3: Clinical expert analysis for Clinical report summarization. R# represents each rubric and C# represents each clinical expert.
-Internal Validation (N=10)
-GPT LLaMa2-7B RO-LMM-S
-Experts R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total R1 / R2 / R3 / R4 / Total $r$
-C1 3.0 / 0.0 / 4.0 / 7.0 / 13.3 2.0 / 0.0 / 0.0 / 7.0 / 9.0 15.0 / 10.0 / 10.0 / 10.0 / 45.0 0.884
-C2 3.0 / 0.0 / 5.0 / 7.0 / 14.4 1.0 / 0.0 / 0.0 / 9.0 / 10.0 15.0 / 10.0 / 9.0 / 10.0 / 44.0
-External Validation (N=10) $r$
-C1 5.0 / 1.0 / 9.0 / 7.0 / 22.0 7.0 / 0.0 / 0.0 / 7.0 / 14.0 10.0 / 9.0 / 8.0 / 8.0 / 35.0 0.977
-C2 6.0 / 3.0 / 9.0 / 7.7 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
-Table S4: Example case for the clinical report summarization task. The red scripts indicate the incorrect or redundant information, and inconsistent format compared to label.
-Clinical Report Summarization Example
-User
-Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
-####\#\## # Input: [MRI report][Pathology report]
-Ground truth
-##\## 1. Breast, RLO 08:00 4cm FN, RLM 04:00 2cm FN, IDC, two, pT1cN1aM0, stage IIA
-- Histologic type: Invasive ductal carcinoma, NOS
-- Tumor margin: Infiltrative
-- Maximum diameter of invasive tumor: 1.5cm and 0.4cm
-- Maximum diameter of invasive and in situ carcinoma: 3.0cm
-- Nuclear grade: 2 - Histologic grade: II (3-2-1)
-- LVI (-) PNI (-) - Intraductal carcinoma component
-[…]
-- Type: Cribriform - Nuclear grade: Low
-- Necrosis: Absent - Location: Periphery
-- Calcification: Present (Malignant abd benign)
-- Associated benign lesion: Intraductal papilloma, sclerosing adenosis and columnar cell lesion
-- RM (-) - Overlying skin (superficial margin): Free of ca (safety margin: 1.4cm)
-- Underlying fascia (deep margin): Free of ca (safety margin: 2.3cm)
-Lymph nodes, total (2/13): PNE (-, Maximal diameter: 3mm)
-- Axillary (0/10): Free of ca
-- Sentinel lymph node1 (1/1), axillary lymph node1 (1/1), axillary lymph node2 (0/1) and axillary lymph node3 (0/0); total (2/3): PNE (-, Maximal diameter: 3mm)
-[…]
-Oncotype Dx: RS 10 CTx omission
-Ours
-##\## 1. Breast, RLO 09:00 4cm FN RMC 03:00 3cm FN, IDC, Multifocal, pT1cN1aM0, stage IIA
-- Histologic type: Invasive ductal carcinoma, NOS
-- Tumor margin: Infiltrative
-- Maximum diameter of invasive tumor: 1.5cm and 0.4cm
-- Maximum diameter of invasive and in situ carcinoma: 3.0cm
-- Nuclear grade: 2 - Histologic grade: II (3-2-1)
-- Lymphovascular invasion: Absent - Perineural invasion: Absent - Intraductal carcinoma component
-[…]
-- Type: Cribriform - Nuclear grade: Low
-- Necrosis: Absent - Location: Periphery
-- Calcification: Present (Malignant abd benign)
-- Associated benign lesion: Intraductal papilloma, sclerosing adenosis and columnar cell lesion
-- Resection margins - Overlying skin (superficial margin): Free of carcinoma (safety margin: 1.4cm)
-- Underlying fascia (deep margin): Free of carcinoma (safety margin: 2.3cm)
-Lymph nodes, total (2/13): PNE (-), max 3mm
-- Axillary (0/10): Free of carcinoma
-- Sentinel lymph node1 (1/1), axillary lymph node1 (1/1), axillary lymph node2 (0/1) and axillary lymph node3 (0/0); total (2/3): Metastatic carcinoma in 2 out of 3 lymph nodes without perinodal soft tissue extension (Maximal diameter: 3mm)
-[…]
-Oncotype DX RS: 20 Adj. CTx omission
-Chat-GPT
-##\##1. Breast, RLO 08:00, 15mm, IDC, pT1cN1(f)M0, Stage IIA ypT2 0cN0M0
-- two (1.5cm and 0.4cm), H2N2
-- two of 13 +
-- LVI (-), DLI (-), HD (-)
-- Extensive IC: (-) - Calcification (+) - RM (-)
-- Overlying skin: Free of ca (safety margin: 1.4cm)
-- Underlying fascia: Free of ca (safety margin: 2.3cm)
-[…]
-Table S5: Example case for the plan suggestion task compared to baseline methods. The red scripts indicate the incorrect answers.
-Example 1
-User
-Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
-##\## Input: [Clinical notes]
-Ground truth
-RT to WB and tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx.
-Ours
-RT to whole breast + tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx.
-Med-Alpaca
-…Breast, RUC, 12:00, 5cm FN, IDC, pT1cN0M0, stage IA…
-ChatDoctor
-… Depending on the patient’s individual circumstances, the radiation oncologist may recommend a course of external beam radiation therapy, brachytherapy, or a combination of the two …
-Chat-GPT
-RT to right breast Conventional fractionation Total dose: 50 Gy in 25 fractions.
-Example 2
-User
-Plan a radiotherapy treatment based on the insights from the radiation oncologist’s clinical note.
-##\##Input: [Clinical notes]
-Ground truth
-RT to Rt. chest wall + AXL Lv 3, IMN, SCL 40.05 Gy / 15 fx.
-Ours
-RT to Rt. chest wall, AXL Lv 3, IMN, SCL 40.05 Gy / 15 fx.
-Med-Alpaca
-.. Status post core biopsy of the right breast, performed …
-ChatDoctor
-… Deliver a dose of 40 Gy in 10 fractions over 4 weeks to the primary tumor, with a margin of 1.1 cm for the overlying skin and 1.2 cm for the underlying fascia. …
-Chat-GPT
-RT to chest wall, IMN, and AXL Conventional fractionation 40.05 Gy / 15 fx.
-Example 3
-User
-Considering the clinical insights from the radiation oncologist’s note, set a treatment plan in terms of radiotherapy.
-##\## Input: [Clinical notes]
-Ground truth
-RT to Lt. WB and tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy in daily fx).
-Ours
-RT to Lt. WB and tumor cavity (SIB) 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy daily fx)
-Med-Alpaca
-.. Lymph nodes, total (0/4): Free of carcinoma - Sentinel lymph node1 (0/1), sentinel lymph node2 (0/1), sentinel lymph node3 (0/1), and axillary lymph node1 (0/1); total (0/4): Free of carcinoma …
-ChatDoctor
-… Radiotherapy will be administered to the breast, left, 2 o’clock, 4 cm FN IDC, pT1bN0M0, H1N2, PNI (-), LVI (-), EIC (-), Calcification: …
-Chat-GPT
-RT to Left breast Conventional fractionation 40.05 Gy / 15 fx.
-Example 4
-User
-Considering the clinical insights from the radiation oncologist’s note, set a treatment plan in terms of radiotherapy.
-##\## Input: [Clinical notes]
-Ground truth
-RT to Lt. WB/Cavity, AXL 2, 3, Lv 4 (ESTRO), IMN 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy in daily fx).
-Ours
-RT to Lt. WB/Cavity, AXL 2, 3, Lv 4 (ESTRO), IMN 40.05 Gy / 48 Gy in 15 fx (2.67 Gy / 3.2 Gy daily fx).
-Med-Alpaca
-[…]
-ChatDoctor
-… These findings suggest that the patient has an aggressive form of breast cancer and may benefit from additional chemotherapy and targeted therapy. …
-Chat-GPT
-RT to left chest wall and axilla Conventional fractionation 40.05 Gy / 15 fx for chest wall 40.05 Gy / 15 fx for axilla.
-Appendix 0.I Examples of Entire Workflow
-We present examples of the entire workflow as indicated in Table S6 and Table S7.
-RO-LMM can take input as a pair of MRI report and Pathology report or another pair of Ultrasound report and Pathology report.
-Our model seamlessly serves as an assistant for radiation oncologists by conducting report summarization, treatment plan suggestion, and 3D Target Volume Segmentation.
-Table S6: Example of Entire Workflow in the case of taking MRI report and Pathology report as input.
-Case 1. MRI report and Pathology report
-User
-Using the provided MRI and pathology findings, design a radiation oncologist’s clinical overview.
-####\#\## # Input: [MRI report][Pathology report]
-RO-LMM-S+
-##\##1. Breast, ROC, 09:00, 3cm FN, IDC, icT2N1(f)M0
-stage IIB (anatomic), stage IIIB (prognostic) ypT0N0M0
-[ic]
-T: - 39mm size mass, H3N3 - LVI (-)
-N: - Multiple suspicious lymph nodes at Rt. axilla level 1, 2, 3 (FNA+)
-[…]
-[yp]
-- No residual ca
-- Fibrosis with hemosiderin laden macrophages and foam cell collection, favoring regressed tumor bed
-- Total LNs (0/25) : Axillary (0/11), Level II (0/6) : SLN (0/4), axillary (non SLN) (0/4)
-- Calcification (+, Benign)
-- RM (-): Nipple, upper, and lower margin: Free of carcinoma (See SF21-04688)
-* Lymphedema prediction
-[…]
-[…]
-[…]
-User
-Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
-####\#\## # Input:[Generated Clinical Note]
-RO-LMM-P++
-RT to Rt. whole breast + AXL Lv 3 + SCL 48 (SIB) - 40.05 Gy / 15 fx.
-User
-Draw the radiotherapy target volume of this patient with following planning CT scan.
-####\#\## # Input:[Generated Treatment Plan][3D CT Scan] patient001.dcm
-RO-LMM-SEG++
-patient001__\__segmentation.dcm
-Table S7: Example of Entire Workflow in the case of taking Ultrasound report and Pathology report as input.
-Case 2. UltraSound report and Pathology report
-User
-Summarize a clinical summary note as a radiation oncologist might, using the provided Ultrasound and pathology details.
-####\#\## # Input: [Ultrasound reports ][Pathology reports]
-RO-LMM-S+
-##\##1. Breast, LUO 02:00 2cm FN, IDC, pT1cN0M0, stage IA
-- Histologic type: Invasive ductal carcinoma
-- Tumor margin: Infiltrative - Nuclear grade: 2 - Histologic grade: II (3-2-1)
-- Maximum diameter of invasive tumor: 1.4cm
-- Maximum diameter of invasive and in situ carcinoma: 2cm
-- Lymphovascular invasion: Not identified - Perineural invasion: Not identified
-[…]
-- Type: Solid / Cribriform - Nuclear grade: Intermediate - Necrosis: Present (Comedo)
-- Location: Central and periphery - Calcification: Present (Malignant and benign)
-[…]
-- Resection margin: - Upper, lower, inner and outer margin: Free of ca
-- Total LNs (0/11): SLN (0/6), non SLN (0/5)
-ER: Positive (AS 8 = PS 5 + IS 3)
-PR: Positive (AS 5 = PS 3 + IS 2)
-HER2: Negative (1+)
-AR: Positive (AS 7 = PS 5 + IS 2)
-[…]
-[…]
-User
-Suggest a radiotherapy approach using the guidance from the given clinical summary of the radiation oncologist.
-####\#\## # Input:[Generated Clinical Note]
-RO-LMM-P++
-RT to Lt. whole breast 26 Gy / 5 fx + 10 Gy / 5 fx boost.
-User
-Draw the radiotherapy target volume of this patient with following planning CT scan.
-####\#\## # Input:[Generated Treatment Plan][3D CT Scan] patient002.dcm
-RO-LMM-SEG++
-patient002__\__segmentation.dcm
-Appendix 0.J Prompt for Evaluation of Treatment Plan Suggestion
-In this paper, we design prompts for evaluating generated treatment plans, including a reference-guided score rubric curated by clinical expert, as shown in Table S9.
-Similar to previous GPT-based evaluation methods, we tailor the prompts combining the specific knowledge of clinical experts.
-By using this prompt, we evaluate the results of our RO-LMM and all the baseline methods using GPT-4.0.
-Table S8: Used prompts with proposed expertise-based rubrics for assessing the performance of treatment plan suggestion.
-Based on <Ground Truth Plan>, score 1 if <Model Generated Plan> satisfies each rubric, else score 0.
-[Rubric]
-R1. Laterality.
-Whether orientation (Rt./Lt.) is correct
-R2. Aim of Surgery.
+Appendix A Details of Training Dataset
+For data preparation, we gathered internal training and validation data from patients treated at Yonsei Cancer Center between January 2009 and December 2022.
+For external validation, data was collected from patients at Yongin Severance Hospital from January 2018 to December 2022, ensuring no overlap with the training set.
+For pre-processing CT scans, all the chest CT scans and target volumes were re-sampled with identical voxel spacing.
+The intensity values of the CT scans were truncated between -1,000 and 1,000 of Hounsfield unit (HU), and linearly normalized to the range of 0 to 1.0.
+During training, 3D patches was randomly cropped, whereas, during inference, the entire 3D CT scan was tested using sliding windows of the same spatial dimension as the training patches.
+Appendix B Details of Open-sourced Dataset
+To rigorously evaluate our framework across different hospital settings, we created synthetic patient cases modeled after Gachon University Gil Hospital.
+These cases covered various breast cancer types (left, right, bilateral), stages (Tis-T3, N0-N3), and included scenarios with breast-conserving surgery, both with and without neoadjuvant chemotherapy, and diverse molecular subtypes.
+GPT-4 was used to generate clinical notes, preoperative ultrasound, MRI reports, and surgical pathology findings tailored to Gil Hospital’s style.
+From 50 generated cases, we selected the 20 most realistic ones with the help of a board-certified radiation oncologist.
+This curated dataset, free of real patient data, is available as an open-source resource at https://github.com/tvseg/ro-lmm.
+Appendix C Detailed Prompt and Rubric for Evaluation
+We design prompts for evaluating generated treatment plans using a reference-guided scoring rubric curated by clinical experts, as detailed in Table A1.
+Our approach builds on previous GPT-based evaluation methods and integrates clinical experts’ knowledge into the prompts.
+Using these tailored prompts, we assess the performance of our RO-LMM model and all baseline methods with GPT-4.0.
+TABLE A1: Score rubrics for treatment plan suggestion.
+[Rubric (R)]
+R1. Laterality. Whether orientation (Rt./Lt.) is correct
+R2.
+Aim of Surgery.
 Whether aim of RT (definitive RT = RT after breast conservation surgery or partial mastectomy.
@@
 In this scenario, chest wall should be irradiated) is correct.
-R3. Treatment Scope.
+R3.
+Treatment Scope.
 Whether detailed extent of radiation field (inclusion or exclusion of WB, CW, AXL, SCL, IMN) and detailed level of inclusion (RTOG = in this guideline, upper SCL node is included / ESTRO = in this guideline, upper SCL is not included) is correct.
-R4. Dose Scheme.
-Whether RT scheme (dose in Gy, and fractions in fx) is correct. It also included SIB or sequential boost dose.
-R5. Hallucination.
-Whether unnecessary information is added or not (If unnecessary & incorrect information was added, score 0 points, otherwise score 1 point)
-[Abbreviations]
-RT.
-Radiotherapy
-BCS.
-Breast Conserving Surgery
-PMRT.
-Post Mastectomy Radiation Therapy
-IMRT.
-Intensity Modulated Radiation Therapy
-PBI.
-Partial Breast Irradiation
-SIB.
-Simultaneous Integrated Boost technique, same as tumor bed boost (sequential), fast-forward protocol
-SCL.
-Supraclavicular Lymph Node
-IMN.
-Internal Mammary Lymph node
-AXL.
-Axilla Lymph Node
-RTOG.
-Radiation Therapy Oncology Group
-ESTRO.
-European Socity for Radiotherapy and Oncology
-CW.
-Chest Wall
-WB.
-Whole Breast
-[Ground Truth Answer Begin]
-<Ground Truth Plan>
-[Ground Truth Answer End]
-[Model Responses Begin]
-<Model Generated Plan>
-[Model Responses End]
-[Answer Format]
-Model, Answer for R1, Answer for R2, Answer for R3, Answer for R4, Answer for R5, (scores ONLY for each Rubric
-in a tabular format, without explanation):
-Table S9: The proposed expertise-based rubrics for assessing the performance of clinical report summarization.
-Based on rubrics, score each clinical report summarization.
-[Rubric]
-Rubric 1: Relevance
-Assessment of whether the necessary information from the Ground Truth (GT) is included, and identification of any extraneous information
-Good (2): The information from the GT is appropriately included.
-Partially correct (1): Some relevant information is included, but there are omissions or irrelevant details.
-Mostly irrelevant (0): The majority of the information is unnecessary or irrelevant.
-Rubric 2: Formatting
-Evaluation of whether the necessary information is presented in the correct format and location:
-Incorrectly placed (0): Information is not presented in the appropriate format or location.
-Correctly placed (1): The necessary information is accurately positioned in the expected format.
-Rubric 3: Length
-Assessment of whether the writing is excessively verbose:
-Too verbose (0): The writing is overly lengthy and lacks conciseness.
-Sufficiently summarized (1): The content is effectively summarized without unnecessary verbosity.
-Rubric 4: Consistency
-Identification of conflicting information between sources:
-Present (0): Conflicting pieces of information are observed.
-Absent (1): There are no conflicting pieces of information between sources.
+R4. Dose Scheme. Whether RT scheme (dose in Gy, and fractions in fx) is correct. It also included SIB or sequential boost dose.
+R5. Hallucination. Whether unnecessary information is added or not (If unnecessary & incorrect information was added, score 0 points, otherwise score 1 point)

### v4

@@
 Inspired by this, here we present RO-LMM, a multi-purpose, comprehensive large multimodal model (LMM) tailored for the field of radiation oncology.
-This model effectively manages a series of tasks within the clinical workflow, including clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation by leveraging the capabilities of LMM.
+This model effectively manages a series of tasks within the clinical workflow, including clinical context summarization, radiotherapy strategy suggestion, and plan-guided target volume segmentation by leveraging the capabilities of LMM.
 In particular, to perform consecutive clinical tasks without error accumulation, we present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the consistency of handling clean inputs.
@@
 Experimental results including multi-centre validation confirm that our RO-LMM with CEFTune and CESEG results in promising performance for multiple clinical tasks with generalization capabilities.
-Index Terms:
-Large multimodal model, Breast cancer, Clinical report, Radiotherapy target volume, Segmentation
-I Introduction
+keywords:
+Large multimodal model , Breast cancer , Clinical report , Radiotherapy target volume , Multimodal segmentation
+journal: Medical Image Analysis
 […]
 figureRO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology.
-The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and 3D target volume segmentation.
-Recently, the emergence of a new generation of AI models inspired by the large language models (LLMs), known as foundation models, marks a significant departure from previous paradigms [1].
+The model seamlessly covers various tasks such as clinical report summarization, radiation radiotherapy strategy suggestion, and 3D target volume segmentation.
+1 Introduction
+Recently, the emergence of a new generation of AI models inspired by large language models (LLM), known as foundation models, marks a significant departure from previous paradigms [36].
 These models are characterized by their massive scale and versatility, which stems from self-supervised training on a diverse array of data.
 These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [3, 11, 13, 28, 33, 26].
-These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision-making.
+These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision making.
 Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [22].
 Specifically, in contrast to these unimodal AIs, generalist medical AIs that synergistically combine foundation models can encompass a holistic understanding of clinical workflows, which can receive a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [43, 41, 51, 36, 47].
-By understanding diverse data types and their interrelationships, multimodal AIs would facilitate more accurate diagnoses, personalized treatment development, and a reduction in medical errors by providing a comprehensive view of patient data.
-In the field of radiation oncology, which the main focus of this article, the integration of multiple modalities holds great importance so that it is one of the most important clinical fields to evaluate the potential of foundation models.
+By understanding various types of data and their interrelationships, multimodal AIs would facilitate more accurate diagnoses, personalized treatment development, and a reduction in medical errors by providing a comprehensive view of patient data.
+In radiation oncology, the primary focus of this article, the integration of multiple modalities is crucial, making it one of the most important clinical fields for evaluating the potential of foundation models.
 Accordingly, here we introduce RO-LMM, a prototype large multimodal model (LMM) specifically designed to support the clinical workflow in radiation oncology.
-In particular, this work significantly extends our related prior work called LLMSeg[13], which focused on multimodal segmentation.
+In particular, this work significantly extends our prior related work called LLMSeg[37], which focused on multimodal segmentation.
 More specifically, RO-LMM enlarges the LLMSeg’s scope by tackling a broader range of clinical tasks in radiation oncology: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes.
-Additionally, it is capable of (2) proposing appropriate treatment plans from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed treatment plans.
-This multifaceted functionality of RO-LLM demonstrates a significant advances toward the expertise of clinical professionals.
-In training LLM to perform these sequential tasks from treatment plan suggestion to target volume segmentation, we found that there exist potentials for error accumulation from each tasks which can potential leads to significant end-to-end performance loss.
+Additionally, it is capable of (2) proposing appropriate radiotherapy strategies from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed radiotherapy strategies.
+This multifaceted functionality of RO-LMM demonstrates significant advances in supporting the expertise of clinical professionals.
+In training LLM to perform these sequential tasks from radiotherapy strategy suggestion to target volume segmentation, we found that there exist potentials for error accumulation from each task which can potential leads to significant end-to-end performance loss.
 Accordingly, another important contribution of this work is the employment and extension of a Noisy Embedding Fine-Tuning (NEFTune) [20] which involves injecting uniform noise into embeddings during the training for each targeted task.
@@
 Furthermore, by expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in novel Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG).
-These advancements prevent from error propagation from each subsequent task and collectively contribute to a marked improvement in the end-to-end model’s generalization capabilities in both internal and external validation.
-As for the proof of the concept study, our RO-LLM framework was adopted for breast cancer study, one of the high-prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging.
+These advancements prevent error propagation between subsequent tasks and collectively contribute to a marked improvement in the end-to-end model’s generalization capabilities in both internal and external validation.
+As for the proof of the concept study, our RO-LMM framework was adopted for breast cancer study, one of the high-prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging.
 Our contributions can be summarized as:
+  • 1.
     We propose a comprehensive framework, dubbed as RO-LMM, wherein LMM assists the broad workflow of radiation oncology for breast cancer treatment.
 To the best of our knowledge, our prototype is the first to assist in the comprehensive workflow of radiation oncology.
-  • To prevent from potential error accumulation throughout consecutive clinical tasks such as clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation, we explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG that significantly enhance the robustness of our method.
-  • Through experiments on various validation settings with real clinical data of breast cancer patients, we demonstrated that RO-LMM outperforms baseline methods.
-II Related Works
-II-A Instruction Fine-tuning in LLM
+  • 2.
+    To prevent from potential error accumulation throughout consecutive clinical tasks such as clinical context summarization, radiotherapy strategy suggestion, and plan-guided target volume segmentation, we explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG that significantly enhance the robustness of our method.
+  • 3.
+    Through experiments on various validation settings with real clinical data of breast cancer patients, we demonstrated that RO-LMM outperforms Default methods.
+Fig.
+1: Schematics of RO-LMM training for three different tasks.
+(a) RO-LMM-S for clinical note summarization.
+(b) RO-LMM-P++ for radiotherapy strategy suggestion.
+(c) RO-LMM-SEG++ for plan-guided target volume segmentation.
+2 Related Works
+2.1 Instruction Fine-tuning in LLM
 In the field of LLM, the instruction fine-tuning has been emerged as a pivotal technique for augmenting model responsiveness.
@@
 Although these models demonstrate robust performance for diverse QA benchmarks, there is a notable gap in their applicability for the practical clinical workflow in a specific domain, such as radiation oncology, which serves as our targeted focus.
-II-B Stabilizing LLM Fine-tuning with Noise
-To enhance the robustness of language models against noisy input, various strategies have been explored.
-Approaches such as SMART [28], FreeLB [29], and R3F [30] employ adversarial training methods, introducing small Gaussian perturbations in embedding dimensions to optimize the model’s performance against noise.
-Aparting from the adversarial training, LSNR [31] takes a different approach by directly optimizing the smoothness of each layer of the language model.
+2.2 Stabilizing LLM Fine-tuning with Noise
+To improve the robustness of language models against noisy input, various strategies have been explored.
+Approaches such as SMART [21], FreeLB [54], and R3F [1], LSNR [18] utilize adversarial training by introducing minor Gaussian perturbations in embedding dimensions, optimizing model performance in noisy environments.
 More recently, NEFTune [20] improves LLM performance during fine-tuning by simply adding random noise into the embedding vectors during training.
@@
 This extension aims to further improve the robustness and generalization capabilities of language models when faced with noisy input.
-Figure 1: Schematics of RO-LMM training for three different tasks.
-(a) RO-LMM-S for clinical note summarization.
-(b) RO-LMM-P++ for radiation treatment plan suggestion.
-(c) RO-LMM-SEG++ for plan-guided target volume segmentation.
-II-C Language-driven Image Segmentation
-Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [32], open-vocabulary segmentation [33, 34], referring segmentation [35], and reasoning segmentation [7].
+2.3 Language-driven Image Segmentation
+Recent research in the field of image segmentation has emerged to incorporate linguistic ability, such like language-driven semantic segmentation [27], open-vocabulary segmentation [12, 31], referring segmentation [49], and reasoning segmentation [26].
 Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable.
@@
 Recent research on clinical context-aware breast cancer radiotherapy delineation [37] has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce.
-By extending [13], we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
-III Methods
+By extending [37], we incorporate radiotherapy strategies generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
+3 Methods
 In this section, we provide a detailed description of our proposed approach designed for sequential text generation tasks, including summarization and suggestions, as well as text-driven image segmentation, whose robustness is improved by consistency embedding finetuning.
 The overall framework is illustrated in Figure 1.
-III-A Consistency Embedding Fine-tuning for Clinical LMM
-To realize the multi-purpose LMM with expertise in clinical report summarization and radiation treatment plan suggestion, we conduct instruction fine-tuning for LLaMA2 [23].
+3.1 Consistency Embedding Fine-tuning for Clinical LMM
+To realize the multi-purpose LMM with expertise in clinical report summarization and radiotherapy strategy suggestion, we conduct instruction fine-tuning for LLaMA2 [46].
 Considering the nuanced differences in the intended objective of each task, we adopt separate training strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert).
@@
 However, for the plan expert, there is a discrepancy between the training and inference scenarios.
-In other words, we use the training set made up of collected summary notes instead of generated notes from the summary expert, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure I.
+In other words, we use the training set made up of collected summary notes instead of generated notes from the summary expert, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding.
 However, at the inference phase, our model takes the generated notes from the trained summary expert.
@@
 More details are as follows.
-Consistency Embedding Fine-Tuning. To enhance the performance of instruction fine-tuning, NEFTune  [14] involves injecting a random noise vector into embeddings during the training process as follows:
+3.2 Consistency Embedding Fine-Tuning.
+To enhance the performance of instruction fine-tuning, NEFTune  [20] involves injecting a random noise vector into embeddings during the training process as follows:
 $\mathcal{L}_{\text{NEFTune}}(\theta)=\mathbb{E}_{{({\mathbf{x}^{\text{emb}}_{\text{txt}}},{\mathbf{y}_{\text{txt}}})\sim D}}\mathcal{L}_{\text{ce}}(f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}),{\mathbf{y}_{\text{txt}}}),$
 ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}={\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\bm{\epsilon,\quad\epsilon}\sim\mathcal{U}(-1,1)$ (1)
-where ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$ is the embedding of data, $B$ denotes batch size, $L$ is token length, $C$ is embedding dimension, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is perturbed embedding, $f_{\theta}(\cdot)$ represents the model parameterized by $\theta$, ${\mathbf{y}_{\text{txt}}}$ is the label of text sample.
+where ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$ is the embedding of data, $B$ denotes batch size, $L$ is token length, $C$ is embedding dimension, $\alpha$ is a tunable parameter.
+The perturbed embedding is denoted as $\tilde{\mathbf{x}}$, and $\epsilon$ represents a random noise vector, drawn from a uniform distribution $\mathcal{U}(-1,1)$.
+Additionally, $\mathbb{E}$ represents the expectation.
+$f_{\theta}(\cdot)$ represents the model parameterized by $\theta$, ${\mathbf{y}_{\text{txt}}}$ is the label of text sample.
 They demonstrated that the effectiveness of incorporating a random noise vector adjusted based on token length, which yields robust results in most fine-tuning scenarios.
-Figure 2: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module.
-These module outputs are aligned through (c) multimodal alignment module.
-To bolster the model’s capability in handling both noisy and clean inputs for treatment plan suggestion task, we incorporate a regularization loss to encourage consistency as follows:
+To bolster the model’s capability in handling both noisy and clean inputs for radiotherapy strategy suggestion task, we incorporate a regularization loss to encourage consistency as follows:
 $\mathcal{L}_{\text{CEFTune}}(\theta)$ $=$ $\mathcal{L}_{\text{NEFTune}}(\theta)+\lambda\mathcal{R}(\theta,\theta^{-}),$ (2)
@@
 Since solely minimizing distance in the same embedding space may weaken the robustness and pose limitations on enforcing semantic similarity, to capture textual similarity between sentences, another contribution of our work is to calculate the distance in the feature space of SentenceBERT [42].
-i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [41].
-III-B LMM-assisted 3D Target Volume Segmentation
+i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using a pretrained model such as SentenceBERT [42].
+Fig.
+2: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module.
+These module outputs are aligned through (c) multimodal alignment module.
+3.3 LMM-assisted 3D Target Volume Segmentation
 For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LLM for context-aware 3D segmentation introduced in LLMSeg[37].
@@
 Additionally, it showed the ability to align target contours with provided clinical text information.
-However, in the extended clinical workflow proposed here, the LMM-assisted segmentation model receives noisy text descriptions of the generated treatment plan from the previous summarization step.
+However, in the extended clinical workflow proposed here, the LMM-assisted segmentation model receives noisy text descriptions of the generated radiotherapy strategy from the previous summarization step.
 To address this noise, we introduce two novel modules: Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG).
@@
 The details are as follows.
-Noise Embedding Segmentation (NESEG).
+3.4 Noise Embedding Segmentation (NESEG).
 Firstly, we introduce the noise augmentation technique, dubbed as NESEG, for the target volume segmentation task.
-We refer to the corresponding RO-LLM as RO-LMM-SEG+.
+We refer to the corresponding RO-LMM as RO-LMM-SEG+.
 Specifically, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, leading to the follow loss function:
@@
 where $\Theta$ = [$\psi,\phi$], ${\mathbf{x}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D CT scan, $B$ denotes batch size, $H$, $W$, and $S$ correspond to height, width, and slice of the CT scan, ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ is the perturbed text embedding, ${\mathbf{y}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D ground truth segmentation mask, $g_{\psi}(\cdot)$ represents the multimodal segmentation model parameterized by $\psi$, $f_{\theta^{*}}$ denotes a frozen LLM, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is the noise injected plan embeddings, and $\mathbf{z}_{\phi}$ is the learnable text prompt parametreized by $\phi$.
-Consistency Embedding Segmentation (CESEG).
+3.5 Consistency Embedding Segmentation (CESEG).
 We further adapt the consistency regularization module, dubbed as CESEG, for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan.
@@
 Apart from the original concept of CEFTune, CESEG modifies the consistency regularization module for the multimodal segmentation task, by combining CEFTune and the text prompt tuning.
-Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is also regularized with respect to the clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$.
+Once the text prompt-prepended noisy embeddings of radiotherapy strategy ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is also regularized with respect to the clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$.
 Specifically, the loss function for CESEG is formulated as:
@@
 We employ the 3D Residual U-Net [8] as for the backbone structure within the image module, as shown in (a).
-In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input.
+In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the radiotherapy strategy embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input.
 In the multimodal alignment module, as illustrated as (c), we adopt both self-attention and cross-attention mechanisms in an interactive manner (image-to-text and text-to-image) for each $l$-th skip-connection layer, by following the concept of promptable segmentation borrowed from Segment Anything Model (SAM) [24]).
@@
 Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module.
-IV Experiments
-IV-A Dataset
-To train our model, we collected internal data cohort composed of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center.
-For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 599 patients with their 3D CT scans and corresponding ground truth masks.
+Table 1: Training data details. CRS: Clinical Report Summarization. RSS: Radiotherapy Strategy Suggestion. PTS: Plan-guided Target Segmentation. US: Ultrasound. Path: Pathology.
+Task Input Output Internal External
+Train Validation Validation
+CRS MRI &\&& US &\&& Path Reports Clinical Note 5,674 121 81
+RSS Clinical Note Radiotherapy Strategy
+PTS Radiotherapy Strategy &\&& CT Scan Segmentation 593 79 81
+4 Experimental settings
+4.1 Dataset
+To train our model, we collected an internal dataset consisting of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center.
+For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 593 patients, including their 3D CT scans and corresponding ground truth masks.
+For the text-related dataset, we established a validation set with 121 patients.
+Due to the limited availability of image data compared to text data, the validation size for the segmentation task was set to 79 patients.
 To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital.
-The detail of dataset is described in Table I and Supplementary Material.
+The detail of dataset is described in Table 1 and A.
 We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals.
 This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
-IV-B Implementation details
-For instruction fine-tuning for RO-LMM-S and RO-LMM-P, we use the LLaMA-2-7B-Chat [23] model as a baseline.
-The maximum context length is set at 4096, and the batch size is set to 2.
-To train RO-LMM-SEG, we employ the 3D Residual U-Net [42] using the open-source library MONAI [^2] https://monai.io/.
-For incorporating the textual information within the multimodel segmentation model, we use the LLaMA-2-7B-Chat model initialized with the pre-trained checkpoint of RO-LMM-P++.
-We set the batch size as 1.
-We use the AdamW [46] optimizer for all the tasks, with a learning rate of 5e-5 until reaching 3 epochs for both RO-LMM-S and RO-LMM-P, and 1e-4 until reaching 100 epochs for RO-LMM-SEG, leveraging 4 NVIDIA A6000 GPUs for each task.
-The hyper-parameter $\lambda$ for CEFTune is set to 1 in all of tasks, and for consistency regularization we adopt the variants of SBERT [41] which is trained on PubMed [^3] https://www.ncbi.nlm.nih.gov/pubmed/ dataset, called PubMedBERT.
-[^4] https://huggingface.co/NeuML/pubmedbert-base-embeddings.
-TABLE I: Training data details. CRS: Clinical Report Summarization. TPS: Treatment Plan Suggestion. PTS: Plan-guided Target Segmentation. US: Ultrasound. Path: Pathology.
-Task Input Output Internal External
-Train Validation Validation
-CRS MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
-TPS Clinical Note Treatment Plan
-PTS Treatment Plan &\&& CT Scan Segmentation 593 79 81
-TABLE II: Quantitative comparison for clinical note summarization. Vanilla: the instruction fine tuning. CI: confidence interval.
-Model Method Internal Validation (N=60) External Validation (N=81)
+Table 2: Quantitative comparison for clinical note summarization. Vanilla: the instruction fine tuning. CI: confidence interval.
+Model Method Internal Validation (N=121) External Validation (N=81)
 Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$ (CI) Rouge L $\uparrow$ (CI) Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$(CI) Rouge L $\uparrow$ (CI)
 Clinical Report Summarization
-LLaMA2 [23] - 0.625 (0.603-0.646) 0.439 (0.423-0.454) 0.625 (0.603-0.646) 0.667 (0.649-0.684) 0.468 (0.456-0.479) 0.667 (0.649-0.684)
-MedAlpaca [25] - 0.529 (0.484-0.578) 0.367 (0.334-0.399) 0.529 (0.484-0.578) 0.414 (0.362-0.462) 0.329 (0.283-0.373) 0.414 (0.362-0.462)
-ChatDoctor [24] - 0.617 (0.572-0.657) 0.411 (0.381-0.439) 0.617 (0.572-0.657) 0.549 (0.508-0.588) 0.425 (0.392-0.459) 0.549 (0.508-0.588)
-Asclepius [27] - 0.662 (0.640-0.682) 0.447 (0.436-0.458) 0.662 (0.640-0.682) 0.590 (0.571-0.609) 0.412 (0.401-0.423) 0.590 (0.571-0.609)
-PMC-LLaMA [26] - 0.398 (0.341-0.451) 0.239 (0.200-0.276) 0.398 (0.341-0.451) 0.382 (0.327-0.434) 0.235 (0.197-0.275) 0.382 (0.327-0.434)
-ChatGPT - 0.672 (0.641-0.703) 0.467 (0.447-0.486) 0.672 (0.641-0.703) 0.651 (0.615-0.683) 0.456 (0.434-0.476) 0.651 (0.615-0.683)
-RO-LMM-S Vanila 0.798 (0.777-0.818) 0.616 (0.598-0.635) 0.798 (0.777-0.818) 0.788 (0.769-0.806) 0.607 (0.586-0.628) 0.788 (0.769-0.806)
-Therapy Plan Suggestion
-LLaMA2 [23] - 0.077 (0.071-0.083) 0.038 (0.034-0.042) 0.077 (0.071-0.083) 0.069 (0.065-0.073) 0.029 (0.027-0.031) 0.069 (0.065-0.073)
-MedAlpaca [25] - 0.139 (0.126-0.152) 0.053 (0.049-0.058) 0.139 (0.126-0.152) 0.137 (0.123-0.152) 0.048 (0.044-0.053) 0.137 (0.123-0.152)
-ChatDoctor [24] - 0.214 (0.194-0.234) 0.082 (0.072-0.091) 0.214 (0.194-0.234) 0.191 (0.178-0.205) 0.065 (0.060-0.071) 0.191 (0.178-0.205)
-Asclepius [27] - 0.176 (0.162-0.190) 0.075 (0.067-0.083) 0.176 (0.162-0.190) 0.142 (0.134-0.149) 0.055 (0.051-0.059) 0.142 (0.134-0.149)
-PMC-LLaMA [26] - 0.118 (0.101-0.137) 0.034 (0.029-0.040) 0.118 (0.101-0.137) 0.165 (0.136-0.195) 0.050 (0.040-0.060) 0.165 (0.136-0.195)
-GPT4.0 - 0.366 (0.343-0.391) 0.230 (0.207-0.252) 0.366 (0.343-0.391) 0.316 (0.294-0.339) 0.249 (0.226-0.272) 0.316 (0.294-0.339)
-RO-LMM-P Vanila 0.651 (0.617-0.689) 0.486 (0.437-0.539) 0.651 (0.617-0.689) 0.591 (0.574-0.609) 0.426 (0.403-0.452) 0.591 (0.574-0.609)
-RO-LMM-P+ NEFT 0.639 (0.607-0.673) 0.467 (0.419-0.516) 0.639 (0.607-0.673) 0.588 (0.573-0.606) 0.419 (0.399-0.441) 0.588 (0.573-0.606)
-RO-LMM-P++ CEFT 0.668 (0.632-0.705) 0.512 (0.460-0.564) 0.668 (0.632-0.705) 0.615 (0.592-0.638) 0.459 (0.429-0.489) 0.615 (0.592-0.638)
-TABLE III: Clinical expert analysis for report summarization. R#: each rubric, C#:each clinical expert.
+LLaMA2 [46] - 0.621 (0.601-0.642) 0.434 (0.419-0.449) 0.621 (0.601-0.642) 0.667 (0.649-0.684) 0.468 (0.456-0.479) 0.667 (0.649-0.684)
+MedAlpaca [15] - 0.514 (0.460-0.549) 0.351 (0.318-0.392) 0.514 (0.460-0.549) 0.414 (0.362-0.462) 0.329 (0.283-0.373) 0.414 (0.362-0.462)
+ChatDoctor [29] - 0.601 (0.544-0.641) 0.400 (0.370-0.424) 0.601 (0.544-0.641) 0.549 (0.508-0.588) 0.425 (0.392-0.459) 0.549 (0.508-0.588)
+Asclepius [25] - 0.652 (0.624-0.674) 0.442 (0.429-0.452) 0.652 (0.624-0.674) 0.590 (0.571-0.609) 0.412 (0.401-0.423) 0.590 (0.571-0.609)
+PMC-LLaMA [50] - 0.390 (0.334-0.436) 0.233 (0.194-0.270) 0.390 (0.334-0.436) 0.382 (0.327-0.434) 0.235 (0.197-0.275) 0.382 (0.327-0.434)
+ChatGPT - 0.657 (0.614-0.685) 0.456 (0.434-0.473) 0.657 (0.614-0.685) 0.651 (0.615-0.683) 0.456 (0.434-0.476) 0.651 (0.615-0.683)
+RO-LMM-S Default 0.783 (0.758-0.806) 0.605 (0.582-0.625) 0.783 (0.758-0.806) 0.788 (0.769-0.806) 0.607 (0.586-0.628) 0.788 (0.769-0.806)
+Radiotherapy Strategy Suggestion
+LLaMA2 [46] - 0.075 (0.068-0.082) 0.036 (0.032-0.040) 0.075 (0.068-0.082) 0.069 (0.065-0.073) 0.029 (0.027-0.031) 0.069 (0.065-0.073)
+MedAlpaca [15] - 0.133 (0.120-0.145) 0.051 (0.047-0.055) 0.133 (0.120-0.145) 0.137 (0.123-0.152) 0.048 (0.044-0.053) 0.137 (0.123-0.152)
+ChatDoctor [29] - 0.205 (0.184-0.225) 0.079 (0.069-0.087) 0.205 (0.184-0.225) 0.191 (0.178-0.205) 0.065 (0.060-0.071) 0.191 (0.178-0.205)
+Asclepius [25] - 0.169 (0.155-0.181) 0.072 (0.064-0.079) 0.169 (0.155-0.181) 0.142 (0.134-0.149) 0.055 (0.051-0.059) 0.142 (0.134-0.149)
+PMC-LLaMA [50] - 0.113 (0.096-0.131) 0.033 (0.028-0.038) 0.113 (0.096-0.131) 0.165 (0.136-0.195) 0.050 (0.040-0.060) 0.165 (0.136-0.195)
+GPT4.0 - 0.356 (0.326-0.375) 0.222 (0.197-0.240) 0.356 (0.326-0.375) 0.316 (0.294-0.339) 0.249 (0.226-0.272) 0.316 (0.294-0.339)
+RO-LMM-P Default 0.634 (0.590-0.670) 0.471 (0.418-0.521) 0.634 (0.590-0.670) 0.591 (0.574-0.609) 0.426 (0.403-0.452) 0.591 (0.574-0.609)
+RO-LMM-P+ NEFT 0.623 (0.576-0.660) 0.453 (0.399-0.510) 0.623 (0.576-0.660) 0.588 (0.573-0.606) 0.419 (0.399-0.441) 0.588 (0.573-0.606)
+RO-LMM-P++ CEFT 0.655 (0.616-0.693) 0.500 (0.449-0.554) 0.655 (0.616-0.693) 0.615 (0.592-0.638) 0.459 (0.429-0.489) 0.615 (0.592-0.638)
+Table 3: Clinical expert analysis for report summarization. R#: each rubric, C#:each clinical expert.
 Internal Validation (N=10)
@@
 C1 3.0 / 0.0 / 4.0 / 7.0 / 13.3 2.0 / 0.0 / 0.0 / 7.0 / 9.0 15.0 / 10.0 / 10.0 / 10.0 / 45.0 0.884
-C2 3.0 / 0.0 / 5.0 / 7.0 / 14.4 1.0 / 0.0 / 0.0 / 9.0 / 10.0 15.0 / 10.0 / 9.0 / 10.0 / 44.0
+C2 3.0 / 0.0 / 5.0 / 7.7 / 15.4 1.0 / 0.0 / 0.0 / 9.0 / 10.0 15.0 / 10.0 / 9.0 / 10.0 / 44.0
 External Validation (N=10) $r$
 C1 5.0 / 1.0 / 9.0 / 7.0 / 22.0 7.0 / 0.0 / 0.0 / 7.0 / 14.0 10.0 / 9.0 / 8.0 / 8.0 / 35.0 0.977
-C2 6.0 / 3.0 / 9.0 / 7.7 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
-IV-C Evaluation Metric and Baseline Models
-The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [47].
-However, acknowledging potential limitations in domain specificity for both tasks, we conduct evaluations using two clinical experts and GPT-4.0.
-A board-certified expert creates five scoring rubrics for plan suggestions, detailed in Table A1.
-Using these rubrics, the experts score each aspect of the plans, assigning 0 or 1 for each rubric separately.
-Similarly, for clinical report summarization, rubrics are created and the same evaluation process is followed.
-Inspired by the recent evaluation methodologies such as GPT-Score [48] and G-Eval [49], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning on the treatment plan task.
-To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [50] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
-To further validate effectiveness of our experiment, we provide the confidence intervals (CI) on each metric by utilizing the non-parametric bootstrap method.
-We calculate CI by conducting 1,000 times of random sampling procedure with replacement from the experimental results, each time equal to the total size of the sample dataset.
-Subsequently, we estimate the 95th percentile CI from the relative frequency distribution of each experimental trial.
-For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [25], ChatDoctor [24], Asclepius [27], and PMC-LLAMA [26].
-Additionally, we include a comparison with ChatGPT [16] for the report summarization task and GPT-4.0 for the treatment plan suggestion task using few-shot in-context learning.
-For the segmentation task, we utilize 3D Residual U-Net [42] and ConTEXTualNET [37] as our baseline methods.
-We further utilize an additional baseline as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
-TABLE IV: Clinical expert analysis for treatment plan suggestion. R#: each rubric, C#: each clinical expert.
+C2 6.0 / 3.0 / 9.0 / 7.0 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
+4.2 Implementation details
+For instruction fine-tuning for RO-LMM-S and RO-LMM-P, we use the LLaMA-2-7B-Chat [46] model as a Default.
+We adopt full-finetuning approach for all of our models.
+The maximum context length is set at 4096, and the batch size is set to 2 for each GPU, leveraging 4 NVIDIA A6000 GPUs.
+We use the AdamW [35] optimizer, with a learning rate of 5e-5 until reaching 3 epochs for both RO-LMM-S and RO-LMM-P.
+The hyper-parameter $\lambda$ for CEFTune is set to 1 in all of tasks, and for consistency regularization we adopt the variants of SBERT [42] which is trained on PubMed [^2] https://www.ncbi.nlm.nih.gov/pubmed/ dataset, called PubMedBERT.333www.huggingface.co/NeuML/pubmedbert-base-embeddings.
+To ensure fair comparisons with other approaches, we utilize the Fully Sharded Data Parallel (FSDP) acceleration method, the bf16 (Brain Floating Point) data format, and gradient checkpointing [4] in all experiments.
+For training the segmentation network, we adopt the 3D Residual U-Net [8] as the backbone architecture, implemented using the MONAI open-source library [^4] https://monai.io/.
+During data preprocessing, all chest CT scans and CTV annotations are re-sampled to a uniform voxel spacing of 1.0 $\times$ 1.0 $\times$ 3.0 mm3.
+Image intensities are truncated to a Hounsfield Unit (HU) range of -1,000 to 1,000 and normalized linearly to fall within a scale of 0 to 1.0.
+For network training, 3D patches sized 384 $\times$ 384 $\times$ 128 pixels are randomly cropped to encompass the entire breast, using a batch size of 1 per GPU.
+Training utilizes 4 NVIDIA A6000 GPUs.
+Performance of the backbone is benchmarked against state-of-the-art models such as 3D SegMsamba [52] and 3D UNETR [16].
+Due to memory constraints, input sizes of 384 $\times$ 384 $\times$ 128 pixels and 320 $\times$ 320 $\times$ 64 pixels are used for SegMamba and UNETR, respectively.
+For evaluation, the complete 3D CT volumes are processed using a sliding window approach across all models.
+To incorporate textual information into the RO-LMM-SEG/SEG+/SEG++ frameworks, we initialize the LLaMA-2-7B-Chat model with the pre-trained RO-LMM-P++ checkpoint.
+For training all the segmentation models, optimization is performed with the AdamW optimizer [35], using a learning rate of 1e-4 over 100 epochs.
+Table 4: Clinical expert analysis for radiotherapy strategy suggestion. R#: each rubric, C#: each clinical expert.
 Internal Validation (N=25)
@@
 GPT-4.0 8.8 / 6.9 / 6.2 / 5.8 / 4.2 / 31.9 8.1 / 7.3 / 6.2 / 1.5 / 8.8 / 31.9 8.5 / 8.1 / 5.0 / 1.2 / 9.2 / 31.9 8.5 / 8.1 / 6.5 / 3.8 / 9.2 / 36.2
-TABLE V: Comparison of 3D target volume segmentation performance.
-Model LMM Tuning Internal Validation (N=79) External Validation (N=81)
-Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
-3D Residual U-Net [42] - 0.804 (0.777-0.831) 0.688 (0.653-0.720) 9.018 (7.077-11.170) 0.690 (0.652-0.725) 0.550 (0.511-0.588) 25.057 (22.844-27.070)
-$\text{ConTEXTualNET}^{\dagger}$ [37] - 0.812 (0.783-0.838) 0.698 (0.661-0.731) 7.392 (5.793-9.083) 0.696 (0.663-0.726) 0.551 (0.517-0.583) 21.253 (20.002-22.474)
-LLaMA2-SEG - 0.817 (0.790-0.841) 0.705 (0.670-0.734) 9.101 (7.189-11.129) 0.692 (0.666-0.714) 0.539 (0.514-0.560) 24.087 (22.450-25.789)
-RO-LMM-SEG Vanilla 0.824 (0.799-0.846) 0.713 (0.680-0.742) 5.528 (3.978-7.295) 0.778 (0.743-0.806) 0.655 (0.621-0.686) 17.459 (15.254-19.678)
-RO-LMM-SEG+ NESEG 0.829 (0.805-0.851) 0.719 (0.690-0.749) 4.636 (3.351-6.073) 0.781 (0.741-0.815) 0.664 (0.626-0.699) 29.410 (27.230-31.386)
-RO-LMM-SEG++ CESEG 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
-modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
-TABLE VI: Comparison of 3D target segmentation performance for overall and specific patient types.
-Types Method Internal Validation External Validation
-#Patients Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) #Patients Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
-All Unimodal 79 0.804 (0.777-0.831) 0.688 (0.653-0.720) 9.018 (7.077-11.170) 81 0.690 (0.652-0.725) 0.550 (0.511-0.588) 25.057 (22.844-27.070)
-Multimodal 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
-Gain $\uparrow$ +0.036 +0.046 -4.627 +0.076 +0.092 -9.821
-Typical Unimodal 55 0.842 (0.821-0.861) 0.734 (0.705-0.761) 7.374 (5.447-9.687) 73 0.723 (0.694-0.749) 0.579 (0.546-0.610) 25.082 (22.783-27.380)
-Multimodal 0.877 (0.866-0.888) 0.784 (0.766-0.802) 3.200 (2.361-4.303) 0.782 (0.742-0.813) 0.659 (0.621-0.690) 14.716 (12.407-17.204)
-Gain $\uparrow$ +0.035 +0.049 -4.173 +0.059 +0.080 -10.366
-Atypical Unimodal 24 0.720 (0.654-0.784) 0.584 (0.513-0.657) 12.519 (8.553-16.584) 8 0.370 (0.178-0.594) 0.259 (0.106-0.449) 27.138 (22.417-30.807)
-Multimodal 0.760 (0.701-0.807) 0.628 (0.562-0.682) 6.948 (3.820-11.014) 0.585 (0.346-0.753) 0.458 (0.269-0.610) 22.410 (16.377-28.314)
-Gain $\uparrow$ +0.040 +0.044 -5.571 +0.215 +0.199 -4.728
-IV-D Results
-Clinical Report Summarization.
-We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table II.
-Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals.
+4.3 Evaluation Metric and Default Models
+The evaluation of generated clinical report summaries and radiotherapy strategies utilizes common NLP metrics, including Rouge-1 (R-1), Rouge-2 (R-2), and Rouge-L (R-L) [32].
+However, acknowledging potential limitations in domain specificity for both tasks, we conduct evaluations using two clinical experts and GPT-4.0.
+A board-certified expert creates five scoring rubrics for plan suggestions, detailed in Table A2.
+Using these rubrics, the experts score each aspect of the plans, assigning 0 or 1 for each rubric separately.
+Similarly, for clinical report summarization, rubrics are created and the same evaluation process is followed.
+Inspired by the recent evaluation methodologies such as GPT-Score [14] and G-Eval [34], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning on the radiotherapy strategy suggestion task.
+To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [10] in millimeter (mm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
+To further validate effectiveness of our experiment, we provide the confidence intervals (CI) on each metric by utilizing the non-parametric bootstrap method.
+We calculate CI by conducting 1,000 times of random sampling procedure with replacement from the experimental results, each time equal to the total size of the sample dataset.
+Subsequently, we estimate the 95th percentile CI from the relative frequency distribution of each experimental trial.
+For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as Defaults, specifically MedAlpaca [15], ChatDoctor [29], Asclepius [25], and PMC-LLAMA [50].
+Additionally, we include a comparison with ChatGPT [38] for the report summarization task and GPT-4.0 for the radiotherapy strategy suggestion task using few-shot in-context learning.
+For the segmentation task, we utilize 3D Residual U-Net [8] and ConTEXTualNET [19] as our Default methods.
+We further utilize an additional Default as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
+5 Results
+5.1 Clinical Report Summarization.
+We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table 2.
+Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the Defaults, providing consistent margins in all metrics and confidence intervals.
 Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning.
-Moreover, we evaluate the generated summaries using expertise-based rubrics by two clinical experts and compare them to baselines, including ChatGPT and LLaMa-2.
-As shown in Table III, our ROB-LMM-S model significantly outperforms all baselines in both internal and external validations, thanks to its domain-specific knowledge.
-Additionally, Pearson correlation ($r$) analysis reveals strong positive inter-clinician correlations ( $>0.85$ and $>0.95$ for internal and external validation, respectively), confirming the reliability of our rubrics and the clinical relevance of ROB-LMM-S.
-Therefore, our ROB-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
-Treatment Plan Suggestion.
-As indicated in Table II, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes.
-Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines.
+Moreover, we evaluate the generated summaries using expertise-based rubrics by two clinical experts and compare them to Defaults, including ChatGPT and LLaMa-2.
+As shown in Table 3, our RO-LMM-S model significantly outperforms all Defaults in both internal and external validations, thanks to its domain-specific knowledge.
+Additionally, Pearson correlation ($r$) analysis reveals strong positive inter-clinician correlations ( $>0.85$ and $>0.95$ for internal and external validation, respectively), confirming the reliability of our rubrics and the clinical relevance of RO-LMM-S.
+Therefore, our RO-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
+5.2 Radiotherapy Strategy Suggestion.
+As indicated in Table 2, we evaluate the model’s performance on radiotherapy strategy suggestion task based on generated clinical notes.
+Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other Defaults.
 Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach.
 However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
-Furthermore, we comprehensively evaluate the generated plans using our expertise-based rubrics and clinical expert analysis, as detailed in Table IV.
+Furthermore, we comprehensively evaluate the generated plans using our expertise-based rubrics and clinical expert analysis, as detailed in Table 4.
 For both internal and external validation, our RO-LMM variants significantly outperform GPT-4.0 with few-shot in-context learning, as assessed by both GPT and clinical experts, due to their domain-specific knowledge.
@@
 The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
-3D Target Volume Segmentation.
-As reported in Table V, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting.
-This can be also observed in qualitative results.
-As shown in Figure 3(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks.
-Moreover, as shown in Figure 3(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows, and RO-LMM-SEG sometimes under-segments the target volume.
-In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated.
-We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type).
-The comparative results are presented in the Table VI separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG.
-Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively.
-When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%.
-This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
-Inference on Public Dataset.
-To validate the effectiveness of our methods when AI-generated results are reused as input, we evaluated the performance by inputting each model’s report summarization results its plan suggestion task.
-We conducted experiments on a publicly available text-related dataset, as detailed in Table VII.
-Our proposed methods consistently outperformed other specialized medical LLMs and variants of GPT (such as ChatGPT and GPT-4) in both clinical report summarization and therapy plan suggestion tasks.
-Notably, our RO-LMM-P++ method demonstrated substantial performance enhancements, with improvements of up to 0.55 and 0.31 over other medical LLMs and GPT variants, respectively, demonstrating our model’s robustness to error accumulation during sequential generations.
-TABLE VII: Quantitative comparison results for our RO-LMM’s clinical report summarization and therapy plan suggestion performance on the publicly available dataset.
+Fig. 3: Qualitative comparison on 3D target volume segmentation task. Red arrows indicate errors.
+Table 5: Comparison of 3D target volume segmentation performance.
+Model LMM Tuning Internal Validation (N=79) External Validation (N=81)
+Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
+3D Residual U-Net [8] - 0.802 (0.775-0.826) 0.517 (0.485-0.545) 0.684 (0.650-0.714) 45.852 (32.441-61.029) 0.689 (0.652-0.723) 0.409 (0.381-0.436) 0.548 (0.510-0.584) 132.825 (113.296-151.525)
+3D SegMamba [52] - 0.576 (0.554-0.598) 0.364 (0.344-0.386) 0.411 (0.389-0.434) 146.610 (140.391-152.103) 0.569 (0.540-0.595) 0.315 (0.295-0.334) 0.408 (0.381-0.434) 156.767 (152.551-160.849)
+3D UNETR [16] - 0.564 (0.544-0.584) 0.340 (0.319-0.361) 0.398 (0.379-0.419) 147.310 (141.625-152.392) 0.538 (0.523-0.551) 0.302 (0.287-0.316) 0.370 (0.357-0.382) 158.272 (154.231-162.241)
+LLMSeg [37] - 0.814 (0.790-0.836) 0.531 (0.499-0.561) 0.699 (0.668-0.727) 42.814 (29.983-56.439) 0.681 (0.656-0.700) 0.397 (0.373-0.420) 0.525 (0.502-0.544) 160.151 (150.742-169.029)
+$\text{ConTEXTualNET}^{\dagger}$ [19] - 0.774 (0.728-0.811) 0.491 (0.451-0.525) 0.664 (0.618-0.703) 31.772 (21.484-44.095) 0.745 (0.708-0.776) 0.444 (0.413-0.472) 0.622 (0.586-0.654) 82.013 (64.103-100.648)
+RO-LMM-SEG Default 0.818 (0.795-0.841) 0.524 (0.493-0.555) 0.705 (0.676-0.736) 20.611 (13.502-29.389) 0.747 (0.715-0.775) 0.419 (0.388-0.449) 0.615 (0.580-0.646) 100.562 (86.656-115.053)
+RO-LMM-SEG+ NESEG 0.828 (0.806-0.850) 0.518 (0.489-0.547) 0.718 (0.690-0.745) 17.805 (10.818-26.427) 0.774 (0.733-0.811) 0.458 (0.427-0.486) 0.659 (0.619-0.695) 34.487 (23.313-46.962)
+RO-LMM-SEG++ CESEG 0.838 (0.818-0.858) 0.555 (0.523-0.583) 0.732 (0.706-0.757) 13.658 (8.133-20.770) 0.761 (0.723-0.794) 0.418 (0.391-0.443) 0.636 (0.601-0.668) 29.459 (21.433-38.557)
+modified for 3D target volume segmentation
+Table 6: Comparison of 3D target segmentation performance for overall and specific patient types.
+Types Method Internal Validation External Validation
+#Patients Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) #Patients Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
+All Unimodal 79 0.802 (0.775-0.826) 0.517 (0.485-0.545) 0.684 (0.650-0.714) 45.852 (32.441-61.029) 81 0.689 (0.652-0.723) 0.409 (0.381-0.436) 0.548 (0.510-0.584) 132.825 (113.296-151.525)
+Multimodal 0.838 (0.818-0.858) 0.555 (0.523-0.583) 0.732 (0.706-0.757) 13.658 (8.133-20.770) 0.761 (0.723-0.794) 0.418 (0.391-0.443) 0.636 (0.601-0.668) 29.459 (21.433-38.557)
+Gain $\uparrow$ +0.036 +0.038 +0.048 -29.194 +0.072 +0.009 +0.088 -103.366
+Typical Unimodal 55 0.837 (0.812-0.860) 0.553 (0.521-0.583) 0.728 (0.694-0.760) 37.449 (21.053-55.604) 73 0.720 (0.691-0.745) 0.429 (0.403-0.454) 0.576 (0.542-0.606) 133.108 (113.319-155.072)
+Multimodal 0.876 (0.865-0.888) 0.603 (0.576-0.630) 0.783 (0.765-0.799) 7.385 (6.095-8.708) 0.776 (0.737-0.808) 0.427 (0.398-0.456) 0.652 (0.615-0.684) 27.519 (19.673-36.031)
+Gain $\uparrow$ +0.039 +0.050 +0.055 -30.064 +0.056 -0.002 +0.076 -105.589
+Atypical Unimodal 24 0.722 (0.658-0.781) 0.435 (0.374-0.494) 0.585 (0.516-0.652) 67.873 (36.956-98.711) 8 0.403 (0.218-0.593) 0.225 (0.108-0.349) 0.292 (0.132-0.460) 128.913 (91.743-157.660)
+Multimodal 0.758 (0.702-0.801) 0.447 (0.394-0.500) 0.624 (0.562-0.674) 28.125 (9.343-50.876) 0.624 (0.441-0.751) 0.328 (0.217-0.425) 0.490 (0.340-0.603) 48.465 (12.673-92.137)
+Gain $\uparrow$ +0.036 +0.012 +0.039 -39.748 +0.221 +0.103 +0.198 -80.448
+Breast-only Unimodal 24 0.837 (0.787-0.878) 0.573 (0.516-0.627) 0.733 (0.668-0.790) 8.040 (4.885-11.685) 51 0.739 (0.705-0.771) 0.452 (0.423-0.481) 0.599 (0.559-0.639) 139.910 (114.241-167.642)
+Multimodal 0.891 (0.870-0.908) 0.625 (0.580-0.669) 0.806 (0.775-0.833) 2.985 (2.365-3.616) 0.786 (0.749-0.820) 0.441 (0.411-0.473) 0.664 (0.625-0.700) 29.320 (20.006-40.850)
+Gain $\uparrow$ +0.054 +0.052 +0.073 -5.055 +0.047 -0.011 +0.065 -110.59
+Table 7: Quantitative comparison results for our RO-LMM’s clinical report summarization and radiotherapy strategy suggestion performance on the publicly available dataset.
 Model Pubilc available dataset (N=15)
-Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$ (CI) Rouge L $\uparrow$ (CI)
+R-1 $\uparrow$ (CI) R-2 $\uparrow$ (CI) R-L $\uparrow$ (CI)
 Clinical Report Summarization
@@
 Ours (RO-LMM-S) 0.795 (0.746-0.833) 0.569 (0.530-0.612) 0.795 (0.746-0.833)
-Therapy Plan Suggestion
+Radiotherapy Strategy Suggestion
 MedAlpaca [15] 0.165 (0.111-0.233) 0.085 (0.064-0.109) 0.165 (0.111-0.233)
@@
 Ours (RO-LMM-P++) 0.669 (0.588-0.740) 0.450 (0.374-0.525) 0.669 (0.588-0.740)
-Figure 3: Qualitative comparison on 3D target volume segmentation task. Red arrows indicate errors.
-TABLE VIII: Ablation study on adopting separate expertise for each textual task against unified strategy.
-Training Strategy Internal (N=60) External (N=81)
-R-1 / R-2 / R-L R-1 / R-2 / R-L
+5.3 3D Target Volume Segmentation.
+As reported in Table 5, we first compare the performance of various 3D segmentation models for target volume segmentation.
+Notably, both 3D SegMamba [52] and 3D UNETR [16] demonstrate inferior performance compared to the 3D Residual U-Net.
+Transformer-based or Mamba-based architectures significantly increase computational memory requirements, which constrains the input image dimensions.
+In the context of radiotherapy segmentation, where the network must process and align volumetric information in the entire volume-level, such input size limitations can be regarded as a significant limitation, especially under constrained computational resources.
+Consequently, we selected the more efficient 3D Residual U-Net as the backbone to maximize the 3D input dimensions and ensure optimal performance.
+Furthermore, our proposed plan-guided 3D target segmentation frameworks (RO-LMM-SEG/SEG+/SEG++) excel other default methods with large gains, specifically remained stable performance across the external validation setting.
+This can be also observed in qualitative results.
+As shown in Figure 3(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks.
+Moreover, as shown in Figure 3(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows, and RO-LMM-SEG sometimes under-segments the target volume.
+In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated.
+Through experiments on consecutive tasks, including radiotherapy strategy suggestion and 3D target volume segmentation, we found that simple noise injection strategies (NEFT and NESEG) are not consistently effective across all tasks.
+In contrast, our proposed CEFT and CESEG methods demonstrate robustness and offer a more generalizable solution within the ROB-LMM framework.
+We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type).
+The comparative results are presented in the Table 6 separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG.
+Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively.
+When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%.
+This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
+Additionally, we specifically categorize cases where regional lymph node treatment is not required and the CTV definition includes only the breast in Table 6, denoted as breast-only, and compare our results with other studies focused on breast segmentation for radiotherapy, as presented in [44].
+Our method demonstrates comparable performance, achieving a Dice score of 0.891 and an HD-95 of 2.985 mm for the internal dataset.
+5.4 Inference on Public Dataset.
+To validate the effectiveness of our methods when AI-generated results are reused as input, we evaluated the performance by inputting each model’s report summarization results its plan suggestion task.
+We conducted experiments on a publicly available text-related dataset, as detailed in Table 7.
+Our proposed methods consistently outperformed other specialized medical LLMs and variants of GPT (such as ChatGPT and GPT-4) in both clinical report summarization and radiotherapy strategy suggestion tasks.
+Notably, our RO-LMM-P++ method demonstrated substantial performance enhancements, with improvements of up to 0.55 and 0.31 over other medical LLMs and GPT variants, respectively, demonstrating our model’s robustness to error accumulation during sequential generations.
+6 Ablation Study
+6.1 Analysis on Strategy of Separate Expert
+To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing another Default model that unifies summarization and plan suggestion, and show the results in Table 8.
+We observe that the unified model’s performance is significantly inferior to our distinct training strategy.
+This implies that summarization task, which provides information within the input, is significantly different from suggesting radiotherapy strategy that does not exist within the input, indicating the challenging nature of the task.
+These results suggest that the knowledge required for each task necessitates task-aware training strategies.
+Table 8: Ablation study on adopting separate expertise for each textual task against unified strategy.
+Training Strategy Internal External
+R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$ R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$
 Clinical Note Summary
-Unified strategy 0.516 / 0.357 / 0.515 0.327 / 0.215 / 0.326
-Distinct strategy 0.823 / 0.635 /0.822 0.805/ 0.634 / 0.804
-Therapy Plan Suggestion
-Unified strategy 0.510 / 0.265 / 0.502 0.409 / 0.197 / 0.406
-Distinct strategy 0.652 / 0.491 / 0.649 0.606 / 0.450 / 0.606
-TABLE IX: Ablation study on CESEG for target segmentation performance with input text variation.
+Unified strategy 0.516 0.357 0.515 0.327 0.215 0.326
+Distinct strategy 0.783 0.605 0.783 0.788 0.607 0.788
+Radiotherapy Strategy Suggestion
+Unified strategy 0.510 0.265 0.502 0.409 0.197 0.406
+Distinct strategy 0.655 0.500 0.655 0.615 0.459 0.615
+Table 9: Ablation study on CESEG for target segmentation performance with input text variation.
 CESEG Input Text Target Segmentation
-Dice $\uparrow$ IoU $\uparrow$
-Ground Truth $0.860_{\pm 0.073}$ $0.761_{\pm 0.103}$
-Generated $0.825_{\pm 0.101}$ $0.714_{\pm 0.134}$
-Difference $\downarrow$ $0.035$ $0.047$
-Ground Truth $0.847_{\pm 0.088}$ $0.743_{\pm 0.111}$
-Generated ${0.840_{\pm 0.094}}$ ${0.733_{\pm 0.120}}$
-Difference $\downarrow$ 0.007 0.010
-V Ablation Study
-Analysis on Strategy of Separate Expert To validate the rationale behind adopting separate strategies for each textual task, we conduct experiments by preparing another baseline model that unifies summarization and plan suggestion, and show the results in Table IX.
-We observe that the unified model’s performance is significantly inferior to our distinct training strategy.
-This implies that summarization task, which provides information within the input, is significantly different from suggesting treatment plan that does not exist within the input, indicating the challenging nature of the task.
-These results suggest that the knowledge required for each task necessitates task-aware training strategies.
+Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI)
+Ground Truth 0.843 (0.823-0.863) 0.559 (0.527-0.591) 0.738 (0.711-0.764)
+Generated 0.818 (0.795-0.841) 0.524 (0.493-0.555) 0.705 (0.676-0.736)
+Difference $\downarrow$ $0.025$ $0.035$ $0.033$
+Ground Truth 0.842 (0.819-0.861) 0.558 (0.524-0.589) 0.736 (0.708-0.761)
+Generated 0.838 (0.818-0.858) 0.555 (0.523-0.583) 0.732 (0.706-0.757)
+Difference $\downarrow$ 0.004 0.003 0.004
+6.2 Analysis on Consistency Regularization
 […]
-TABLE X: Inference computational complexity. RM: RO-LMM.
+6.3 Analysis of Noise type and Intensity ($\alpha$)
+Since our proposed method involves injecting noise into embeddings, we conduct a study to analyze the effects of both noise type and intensity under various conditions, as shown in Table 10.
+First, regarding noise type, we observe that the method performed best with uniform noise compared to Gaussian noise, leading us to adopt uniform noise as the default configuration.
+We then examine noise intensity by varying the parameter $\alpha$ from 5 to 15.
+Lower values of $\alpha$ result in improved performance on both internal and external datasets, while excessively high values caused a noticeable drop in performance on the external dataset.
+Building on the main motivation behind NEFTune [20], we hypothesize that introducing noise to the embedding during training reduces overfitting to the specific characteristics of the instruction fine-tuning dataset, such as formatting details and exact wording.
+This allows the model to generate responses that incorporate broader knowledge rather than simply replicating the instruction distribution.
+However, increasing the noise intensity excessively can cause a mismatch between the original data distribution and the test set distribution, underscoring the importance of selecting an optimal noise intensity.
+Based on these empirical results, we selected $\alpha=5$ as the baseline.
+6.4 Analysis of Finetuning Approach and Model Backbone
+In our work, we adopt the full-finetuning approach for instruction finetuning of LLMs.
+To validate this choice, we conduct experiments comparing full-finetuning to parameter-efficient methods, such as LoRA as indicated in Table 10.
+The results showed that using LoRA led to a significant drop in performance on both internal and external datasets.
+Given the specialized nature of our scope, focused on medical professionals, parameter-efficient finetuning proved insufficient for capturing the target domain knowledge.
+These findings indicate that full-finetuning is essential for adapting LLMs to the medical domain, particularly in the context of radiation oncology for breast cancer.
+We conduct experiments using the larger 13B LLaMA-2 backbone version and observe a similar performance trend compared to the 7B model as shown in Table 10.
+Despite the larger backbone size, our proposed CEFT method outperforms the other baselines, demonstrating scalability.
+Taking into account both cost-effectiveness and performance, we adopt the 7B model as the default configuration.
+Table 10: Component analysis of our proposed method on radiotherapy strategy suggestion.
+Component Condition Ours Internal Validation External Validation
+R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$ R-1 $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$
+Noise type Uniform 5 $\checkmark$ 0.655 0.500 0.655 0.615 0.459 0.615
+Gaussian 5 0.625 0.449 0.629 0.588 0.425 0.586
+Noise Intensity ($\alpha$) 5 $\checkmark$ 0.655 0.500 0.655 0.615 0.459 0.615
+10 0.648 0.488 0.647 0.597 0.435 0.598
+15 0.642 0.481 0.642 0.576 0.415 0.572
+Finetuning Full-finetune $\checkmark$ 0.655 0.500 0.655 0.615 0.459 0.615
+LoRA 0.384 0.352 0.373 0.325 0.301 0.323
+Backbone LLaMA-13B 0.645 0.478 0.640 0.597 0.436 0.594
++NEFT 0.658 0.500 0.653 0.601 0.438 0.597
++CEFT $\checkmark$ 0.662 0.502 0.661 0.613 0.441 0.603
+Table 11: Inference computational complexity.
 External Validation (N=81)
-RM-S RM-P++ RM-SEG++ Analysis
+RO-LMM-S RO-LMM-P++ RO-LMM-SEG++ Analysis
 Memory (GB) 13.68 13.57 17.89 Mean 14.93
 Duration (sec/patient) 6.89 $\pm$ 1.68 1.75 $\pm$ 0.51 1.54 $\pm$ 0.72 Total 10.19
-VI Discussion
+7 Discussion
 While our study provides valuable insights into the application of radiation oncology, several limitations must be acknowledged.
@@
 Future research should focus on developing a framework capable of delineating these boost volumes to aid in clinical treatment.
+Thirdly, the inclusion of non-standard dose prescriptions reflecting country- or institution-specific practices in the training data can limit the generalizability of the results and cause confusion for non-experts.
+To ensure safer and more broadly applicable clinical implementation, it is crucial to train models using standardized, evidence-based prescriptions.
 Finally, our model did not utilize all clinical information from the actual EMR for decision-making due to token length limitations.
@@
 Overcoming token length constraints to integrate all EMR data will be essential for future studies.
-Despite limitations, our RO-LMM demonstrates the potential to efficiently handle a broad range of clinical workflows using a single NVIDIA A6000 GPU within a 24GB memory, completing tasks in approximately 10 seconds (see Table X).
+Despite limitations, our RO-LMM demonstrates the potential to efficiently handle a broad range of clinical workflows using a single NVIDIA A6000 GPU within a 24GB memory, completing tasks in approximately 10 seconds (see Table 11).
 This capability facilitates local implementation even with limited GPU resources and holds the promise of reducing the workload of radiation oncologists via the automation of clinical report-based plan suggestion for delineating radiation target volumes.
@@
 These results suggest that our locally tuned RO-LMM-P++ holds significant potential to replace closed-source LLMs while maintaining data privacy and compliance with local data exposure regulations.
-VII Conclusion
+8 Conclusion
 In this work, we introduce RO-LMM, a multi-purpose, comprehensive foundation model tailored for radiation oncology.
-Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks encompassing overall workflow of radiation oncology: clinical report summarization, radiation treatment plan suggestion, and plan-guided 3D target volume segmentation.
+Addressing limitations in current medical AI models confined to specific tasks, RO-LMM demonstrates proficiency in diverse tasks encompassing overall workflow of radiation oncology: clinical report summarization, radiotherapy strategy suggestion, and plan-guided 3D target volume segmentation.
 Another key contribution of this work is the introduction of consistency technique into both text and segmentation task.
@@
 These findings mark a significant stride toward developing a versatile AI model, hinting at the potential for a multi-purpose medical AI model in radiation oncology.
-Ethical Statement All the patient information and any potential personally identifiable information from used datasets are de-identifiied, and clinical data usage is approved by the Institutional Review Board (IRB).
-Appendix A Details of Training Dataset
+Acknowledgments
+This research was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (#2022R1A2C2008623), and also by the Ministry of Education (#RS-2023-00242164).
+This work was also supported by the Korea Health Industry Development Institute (KHIDI) grant funded by the Ministry of Health & Welfare, Republic of Korea (#HI23C0730000023).
+This work was also supported by the National Research Foundation of Korea (NRF) (#RS-2023-00262527) and the Korea Medical Device Development Fund grant funded by the Korea government (the Ministry of Science and ICT, the Ministry of Trade, Industry and Energy, the Ministry of Health & Welfare, the Ministry of Food and Drug Safety) (Project Number: 1711137899, KMDF_PR_20200901_0015).
+Appendix A Details of Training and Validation Dataset
 For data preparation, we gathered internal training and validation data from patients treated at Yonsei Cancer Center between January 2009 and December 2022.
 For external validation, data was collected from patients at Yongin Severance Hospital from January 2018 to December 2022, ensuring no overlap with the training set.
-For pre-processing CT scans, all the chest CT scans and target volumes were re-sampled with identical voxel spacing.
-The intensity values of the CT scans were truncated between -1,000 and 1,000 of Hounsfield unit (HU), and linearly normalized to the range of 0 to 1.0.
+For pre-processing, all the chest CT scans and target volumes were re-sampled with identical voxel spacing.
+The Hounsfield unit (HU) of the CT scans were truncated between -1,000 and 1,000, and linearly normalized to the range of 0 to 1.0.
 During training, 3D patches was randomly cropped, whereas, during inference, the entire 3D CT scan was tested using sliding windows of the same spatial dimension as the training patches.
+All the patient information and any potential personally identifiable information from used datasets are de-identifiied, and clinical data usage is approved by the Institutional Review Board (IRB).
 Appendix B Details of Open-sourced Dataset
@@
 Appendix C Detailed Prompt and Rubric for Evaluation
-We design prompts for evaluating generated treatment plans using a reference-guided scoring rubric curated by clinical experts, as detailed in Table A1.
+We design prompts for evaluating generated report summarizations and radiotherapy strategies using a reference-guided scoring rubric curated by clinical experts, as detailed in Table A1 and  A2, respectively.
+For the report summarizations, there are four evaluation criteria, with the first criterion having a weight of 2, making the total possible score 5.
+For the radiotherapy strategy suggestions, there are five criteria, each contributing a score of 1, resulting in a total score of 5.
 Our approach builds on previous GPT-based evaluation methods and integrates clinical experts’ knowledge into the prompts.
-Using these tailored prompts, we assess the performance of our RO-LMM model and all baseline methods with GPT-4.0.
-TABLE A1: Score rubrics for treatment plan suggestion.
+Using these tailored prompts, we assess the performance of our RO-LMM model and all Default methods with GPT-4.0.
+Table A1: The proposed expertise-based rubrics for assessing the performance of clinical report summarization.
+Based on rubrics, score each clinical report summarization.
+Rubric 1: Relevance Assessment of whether the necessary information from the Ground Truth (GT) is included, and identification of any extraneous information
+Good (2): The information from the GT is appropriately included.
+Partially correct (1): Some relevant information is included, but there are omissions or irrelevant details.
+Mostly irrelevant (0): The majority of the information is unnecessary or irrelevant.
+Rubric 2: Formatting Evaluation of whether the necessary information is presented in the correct format and location:
+Incorrectly placed (0): Information is not presented in the appropriate format or location.
+Correctly placed (1): The necessary information is accurately positioned in the expected format.
+Rubric 3: Length Assessment of whether the writing is excessively verbose:
+Too verbose (0): The writing is overly lengthy and lacks conciseness.
+Sufficiently summarized (1): The content is effectively summarized without unnecessary verbosity.
+Rubric 4: Consistency Identification of conflicting information between sources:
+Present (0): Conflicting pieces of information are observed.
+Absent (1): There are no conflicting pieces of information between sources.
+Table A2: Score rubrics for radiotherapy strategy suggestion.
 [Rubric (R)]
@@
 R5. Hallucination. Whether unnecessary information is added or not (If unnecessary & incorrect information was added, score 0 points, otherwise score 1 point)
+[Abbreviations]
+RT. Radiotherapy
+BCS. Breast Conserving Surgery
+WB. Whole Breast
+CW. Chest Wall
+AXL. Axilla Lymph Node
+SCL. Supraclavicular Lymph Node
+IMN. Internal Mammary Lymph node
+RTOG. Radiation Therapy Oncology Group
+ESTRO. European Socity for Radiotherapy and Oncology
+SIB. Simultaneous Integrated Boost technique, same as tumor bed boost (sequential), fast-forward protocol
+[Ground Truth Answer Begin]
+$<$Ground Truth Plan$>$
+[Ground Truth Answer End]
+[Model Responses Begin]
+$<$Model Generated Plan$>$
+[Model Responses End]
+[Answer Format]
+Model, Answer for R1, Answer for R2, Answer for R3, Answer for R4, Answer
+for R5, scores ONLY for each Rubric in a tabular format, without explanation: